### v1

# Breaking the Trend: How to Avoid Cherry-Picked Signals

Abstract
Our empirical results, illustrated in Fig.5, show an impressive fit with the pretty complex theoritical Sharpe formula of a Trend following strategy depending on the parameter of the signal, which was derived by Grebenkov and Serror (2014). That empirical fit convinces us that a mean-reversion process with only one time scale is enough to model, in a pretty precise way, the reality of the trend-following mechanism at the average scale of CTAs and as a consequence, using only one simple $\mathrm{EMA}$, appears optimal to capture the trend. As a consequence, using a complex basket of different complex indicators as signal, do not seem to be so rational or optimal and exposes to the risk of cherry-picking.
April 15, 2025
I thank Prof Campbell Harvey, Robert Kosowsky, Valeriy Zakamulin, Denis Grebenkov and Raphael Douady for their feedbacks and valuable insights.
1 Introduction
The usual recipe to determine a portfolio in the CTA’s industry is to use a blend of many different technical indicators as signal for each underlying instrument (most often individual indicator gives either a long or a short position). Then, a risk management process is applied to the signals to ensure that the portfolio is diversified enough and target a constant risk. One simple solution is to impose an equal conditional risk in each asset class, as the SG Trend Indicator does and to size the portfolio to target a constant volatility. Another solution is to use the correlation matrix and apply mathematical optimization to maximize the reward/risk ratio. The well-known Markowitz solution set positions as linearly depending on the signals. The linear dependence is derived simply througth the normalization of the signals by the inverse of the correlation matrix between returns, but the results are not appealing. The Agnostic Risk Parity ($\mathrm{ARP}$), introduced by Benichou at al (2017), normalizes the signals through the inverse of the square root of the correlation matrix. Benichou at al (2017) proposed this portfolio because it was rotationally invariant. Since this concept is primarily understood by physicists, their portfolio is not yet widely accepted in the finance community, with only a few citations, despite posing a serious challenge to Markowitz.
In reality, one hidden assumption of Markowitz optimization is that the expected returns from the signals are known and certain, but this assumption is incorrect. Valeyre (2024) proved that the Agnostic Risk Parity ($\mathrm{ARP}$) approach was optimal when the correlation matrix between the trends or signals was a kind of random matrix where a dominant factor of this correlation matrix was very dominant and diffused randomnly. Valeyre (2024) suggested an interacting agents model to justify such an hypothesis.
Most of the popular technical indicators are based on moving averages of past prices. The most popular is the Simple Moving Average ($\mathrm{SMA}$), while less commonly used types include the Linear Moving Average and the Exponential Moving Average on returns ($\mathrm{EMA}$). Each moving average is computed using an averaging window of a particular size. Trend-following indicators can also be based on a combination of moving averages, such as crossovers (one with a short window size and another with a long window size). The switch is determined when the short-term moving average crosses the long-term moving average. As an example, The Societe Generale Trend signal, which is a reference among CTAs, is detemined only by the crossover based on a simple average based 20 and 120 days parameters. Momentum ($MOM$) is also a very popular indicator and is defined as the difference between two prices. Hurst et al (2017) analysed the performance simulated on the last century using a mixture of 1, 3 months and 1 year momentum. Lamperiere et al (2014) analysed the performance simulated on the last 2 centuries using a 50 days EMA signal. The Bollinger Band ($\mathrm{BB}$) is very popular but non continuous with only 3 outputs 0 and 1 or -1 when price is oustide a band. The Moving Average Convergence/Divergence ($\mathrm{MACD}$) uses a combination of three EMAs to capture mean reversion at both short- and long-term scales while identifying trends in the long term.
We can describe some of these indicators through their sensitivity to past daily returns as introduced by Zakamuli and Giner (2020). Fig.1 displays some cases. It can be observed that the Exponential Moving Average ($\mathrm{EMA}$) exhibits a sensitivity that declines exponentially, while the crossover resembles a hat shape. Sensitivity is low for very recent past returns, high for a moderately distant past, and negligible for a long past. The shape of the crossover can be adapted if price behavior exhibits mean-reverting tendencies on very short time scales, as seen in the case of individual stocks with low liquidity. From an academic point of view, almost all these trend-following rules are ad hoc and lack theoretical justification for their optimality. Zakamuli and Giner (2020) related them to each other through the analysis of their sensitivity to past daily returns (so a 33 days EMA should be the optimal EMA to replicate the mixuture of 1, 3 months and 1 year momentum used by Hurst et al (2017) while a 43 days EMA should be the optimal one for the SG trend Indicator crossover). Zakamuli and Giner (2020) also reminds us that it is well known that if returns follow an autoregressive process, the best predictor has the same functional form as the autoregressive process being predicted. Therefore, the expected structure of the autoregressive process should provide a good proxy for the best predictor.
Zakamuli and Giner (2024) shows that the indicator with the same sensitivity to past returns as the autoregressive coefficients of the return process is both the best predictor and the optimal indicator, yielding the highest Sharpe ratio when the investor has only two options (Buy or Sell with the same level of risk). He also employs a two-state regime-switching model (bull and bear regimes), a widely accepted framework for stock returns (Timmermann (2000); Fruhwirth-Schnatter (2006); Giner and Zakamulin (2023)), using a semi-Markov model to detect negative autocorrelations over very long time scales. He argues that, within this framework, the MACD indicator with three different time scales could be adapted to construct the optimal portfolio.
However, mean reversion at larger scales is unconvincing, and the two-state regime does not seem natural for modeling trend-following mechanisms, which are better explained by herding behavior rather than by an external guiding force. Models in which agents partially imitate each other are more likely to produce continuous regime changes rather than abrupt shifts between two states. Consequently, a more realistic approach would require a more complex model than a simple two-state regime. A more credible representation would assume that trends follow an Ornstein-Uhlenbeck process or a similar mean-reverting process.
Grebenkov and Serror (2014) proved that, in the more complex yet more insightful case where portfolio positions depend linearly on the signal, the $\mathrm{EMA}$ with a larger decay parameter than the decay parameter of the autoregressive coefficients of the return process yields the optimal portfolio. This finding is in stark disagreement with Zakamuli and Giner (2024), simply because Grebenkov and Serror (2014) solved a different, yet superior, optimization problem. Indeed, assuming that positions depend linearly on the signal appears much more natural and is not as suboptimal as merely assuming a binary choice between a long or short position with the same level of risk withtout the possibility of implementing ”money management”. Most important the version of Grebenkov and Serror (2014) is in line with Markowitz optimization and Agnostic risk parity by Benichou at al (2017) where positions are linearly depending on the signals.
Moreover, Grebenkov and Serror (2014) is the only one to have derived an elegant and rigorous formula for the theoretical Sharpe ratio, explicitly linking it to the parameters of the return process and the relaxation time of the $\mathrm{EMA}$. That formula is key in our main objective as we want to validate it using empirical measurements.
Two years later, while Grebenkov and Serror (2014) remained largely unrecognized by the academic finance community, the literature continued to focus on the wrong optimization problem—one where positions do not depend linearly on the signals. Dai at al (2016) addressed this issue by imposing a constraint that allowed only two possible positions (long or zero). Similarly, Nguyen at al (2014b) determined the optimal threshold for triggering either a long or short position.
Our first objective was to validate empirically the model of Grebenkov and Serror (2014) through testing their beautiful formula describing the sensitivity of the empirical Sharpe ratio of the ’Agnostic Risk Portfolio’ ($\mathrm{ARP}$)—one of the components of the optimal trend-following portfolio derived by Valeyre (2024)—to the parameter of the $\mathrm{EMA}$. We also wanted to explore more complex signal than $\mathrm{EMA}$ to confirm that one time scale is enough to describe trends. We decided to limit the other signals to a combination of three $\mathrm{EMA}$s to align with the $\mathrm{MACD}$ which could make sense if the autoregressive structure of returns could be more complex than the one which fits perfectly for a $\mathrm{EMA}$. However, there are two major differences:
  • We aim to set the slope of the sensitivity at lag zero to zero as we focuse only on systematic returns through cross asset futures excluding specific risk. We prefer to avoid a positive slope, as it would be curious for recent returns to have less impact on the trend than older ones and as short term Mean reversion behavior is known to be less important than for single stocks.
  • We aim to achieve a larger fat tail for the larger scales instead of having a contrarian contribution which appears more adapted for capturing residual risk in single stocks but not for systematic part of returns.
Our second objective was to understand that a simple Exponential Moving Average ($\mathrm{EMA}$), which is supposed to be both optimal and simple, could also be replicated by a highly complex but more usual combination of indicators, such as Bollinger Bands, whose elementary indicator is even nonlinear and path-dependent. This second objective raises the question: why use a complex combination of complex indicators that are sensitive to cherry-picking when a simple and elegant solution already exists?
(a) MOM 100 days
(b) SMA 100 days
(c) EMA 100 days
(d) SMA Crossover 30-100 days
(e) EMA Crossover 30-100 days
(f) MACD 30-100-400 days
Figure 1: Capture of Zakamuli and Giner (2024). ”Spectrum” of different indicators based on Zakamuli and Giner (2024) i.e. sensistivity to past daily returns. $\mathrm{MOM}$ is the different between two prices, $\mathrm{SMA}$ is the usual moving average on prices, $\mathrm{EMA}$ is the exponential moving avarage of returns, $\mathrm{SMA}$ crossover is the difference between 2 $\mathrm{SMA}$, $\mathrm{EMA}$ crossover is the difference between 2 $\mathrm{EMA}$. $\mathrm{MACD}$ has 3 time scales and could be adapted according Zakamuli and Giner (2024) which is contrarian on the large scales
2 Theory from Grebenkov and Serror (2014)
Grebenkov and Serror (2014), who solved the correct optimization problem—maximizing the Sharpe ratio at the portfolio level while ensuring that positions remain linearly dependent on signals—introduced a diffusive process with a positive autocorrelation model for returns $r_{i,t}$ of the $i$th instrument among the N ones to describe trends in financial market prices , as formulated in Eq.1.
$r_{i,t}=\epsilon_{i,t}+\beta\sum_{k=1}^{t-1}{\left(1-\lambda\right)^{t-1-k}\xi_{i,k}}$ (1)
The weights $\left(1-\lambda\right)^{t-1-k}$ are declining exponentialy. The parameter $\lambda$ describes the inverse of the relaxation time of the autocorrelation, and $\beta$ represents the strength of that autocorrelation component in the returns. The noises $\xi_{i,k}$ are independent in the $k$ axis but may be correlated with $i$ axis (Valeyre (2024) introduces particular cases). The noises $\epsilon_{i,t}$ are independent in the $t$ axis. Through that model, the trend is $\beta\sum_{k=1}^{t-1}{\left(1-\lambda\right)^{t-1-k}\xi_{i,k}}$ follows an auto regressive model of order 1 or an Orstein-Uhlenbeck process if time lag is infinitesimally small which is a mean reversion usual and natural process in complex systems with only one time scale.
Through variogram measurements on the Dow Jones index, they estimated that $\lambda=0.01$ and $\beta_{0}=0.1$ (Fig.2). The fit appears suspiciously perfect, considering the expectation of multiple time scales with long memory among investors. Additionally, the result is not entirely convincing, as the 100-year period may be heterogeneous, and the fit is less accurate at shorter time scales, where measurements should, in principle, be less noisy. Another drawback is that the fit may be specific to the Dow Jones Index, the only index with such a long history. To say it in another way, we can suspect the presence of two distinct time scales, as the fit is not perfect for time scales shorter than 100 days. Another possible explanation is that the autocorrelation may not have been consistent over such a long period.
Figure 2: Capture of Grebenkov and Serror (2014). Variogram applied to Dow Jones (1900-2012)to fit $\beta_{0}=0.08$ and $\lambda=0.011$ parameters
Instead of focusing on the best predictor or on the optimal portfolio with suboptimal constraints—such as the possibility of being either long with the same risk or flat, which is unfortunately common in the literature—Grebenkov and Serror (2014) assumes that positions should better depend linearly on the signals, which is a natural approach which is in agreement with the Markowitz solution. They then determine the optimal signal that maximizes the Sharpe ratio. Thus, Grebenkov and Serror (2014) introduced substantial novelty, even though it is not yet widely considered by most academics specializing in trend-following.
In detail, they derived an explicit formula in Eq.3 for the optimal parameter ($\eta_{\text{opt}}$) of the Exponential Moving Average ($\mathrm{EMA}$) applied to returns as a trend indicator for a trend-following strategy (Eq.6 mathematically defines later this indicator). Additionally, they derived in Eq.4 the theoretical Sharpe ratio function ($\mathrm{SHARPE}\left(\eta\right)$), assuming returns are generated through their autocorrelation model (1), as a function of $\eta$ and $\lambda$, while also considering the inclusion of trading costs, denoted as $\theta$.
$\beta=\beta_{0}\sqrt{\lambda\left(2-\lambda\right)}$ (2)
$\eta_{\text{opt}}=\lambda\sqrt{1+2\frac{\beta^{2}_{0}}{\lambda}}$ (3)
$\mathrm{SHARPE}\left(\eta\right)=\frac{\beta^{2}_{0}\sqrt{2\eta}-\frac{2}{\pi}\theta\sqrt{\eta}\left(\lambda+\eta\right)}{\sqrt{\left(\lambda+\eta\right)^{2}+2\beta^{2}_{0}\left(\lambda+\eta\right)}}$ (4)
It is interesting to observe from Eq.3 that $\frac{1}{\eta_{\text{opt}}}$, the time scale of the signal $\mathrm{EMA}$ generating the optimal strategy is always shorter than $\frac{1}{\lambda}$, the time scale of the best predictor $\mathrm{EMA}$ (or the optimal indicator in the sub optimal optimization problem of Zakamuli and Giner (2024)) and that of the autocorrelated process of returns. Indeed, when $\beta_{0}$ increases (either due to a stronger trend or a more diversified universe), the time scale of the optimal $\mathrm{EMA}$ should be reduced. As a conclusion, theoretically the parameters of the signals should be adjusted when the universe is increased.
Fig.3 illustrates the very complex theoritical formula Eq.4 through a simple graph which enable to undertsand its sensitivity against the parameter of the $\mathrm{EMA}$.
[…]
3 Description of the empirical analysis
3.1 Single $\mathrm{EMA}$ indicator
The $\mathrm{EMA}$ at the time $t$ is a vector of $N$ exponential moving averages of normalized returns, one for each of the $N$ instruments. The $EMA$ of the instrument $i$ at time $t+1$ is defined by Eq.6. The incrementation is on daily basis but it could be generalized to minutes returns.
$\sigma^{2}_{i,t+1}\left(\eta\right)=\left(1-\eta\right)\sigma^{2}_{i,t+1}\left(\eta\right)+\eta r^{2}_{i,t+1}$ (5)
[…]
The indicator $\mathrm{EMA}\left(150\right)$ is defined by the $\mathrm{EMA}_{i,t+1}\left(\eta\right)$ when $\eta=\frac{1}{150}$ is applied to every underlying $i$ and time $t$. We use $\sqrt{\eta}$ in Eq.6 so that the std of $\mathrm{EMA}_{i,t+1}\left(\eta\right)$ is 1 theoritically if we have a random walk.
3.2 $\mathrm{MACD}$ as a 3 time scales $\mathrm{EMA}$ indicator
Inspired by $\mathrm{MACD}$, we introduced a combination of $\mathrm{EMA}$ applying the Eq.7 while determining $\omega_{1}$ so that the derivative of the sensitivity to past daily returns is at 0 at the lag 0 and replicate curves in Fig.4
$\mathrm{MACD}_{i,t}\left(\eta_{1},\eta_{2},\eta_{3},\omega_{1},\omega_{2},\omega_{3}\right)=\omega_{1}\mathrm{EMA}_{i,t}\left(\eta_{1}\right)+\omega_{2}\mathrm{EMA}_{i,t}\left(\eta_{2}\right)+\omega_{3}\mathrm{EMA}_{i,t}\left(\eta_{3}\right)$ (7)
The derivative at zeros yields to Eq.8.
$0=\omega_{1}\sqrt{\eta_{1}}+\omega_{2}\sqrt{\eta_{2}}+\omega_{3}\sqrt{\eta_{3}}$ (8)
3.3 Agnostic Risk Parity ($\mathrm{ARP}$) when smoothing the portfolio and targeting a constant volatility
We first estimated the correlation matrix $C$, of dimension $N\times N$, using a 750-day exponential moving average applied to weekly returns when implemnting the RIE filter introduced by Bun at al (2016). The vector $\Sigma$ consists of N values representing the standard deviations estimated using a 40-day exponential moving average applied to daily returns.
Next we expressed the positions vector as a linear function of the signals ($\mathrm{EMA}$ or $\mathrm{MACD}$) applying a normalisation that involves the inverse of the square root of the estimated correlation matrix and volatilities . This follows the formulation of the agnostic risk parity portfolio (${\mathrm{ARP}}$) introduced by Benichou at al (2017)
We then used the parameter $\rho=\frac{1}{20}$ for portfolio smoothing as specified in Eq.LABEL:eq:arp2.
Finally, wer applied a resizing process in Eq.13 to target a constant volatility for the final positions ${\mathrm{ARP}}$ which is a vector of N weights at time $t+1$.
$\left\{\begin{array}[]{l}\hat{\mathrm{ARP}}_{t+1}=\left(1-\rho\right)\hat{\mathrm{ARP}}_{t}+\rho\Sigma^{-1}_{t}C^{-0.5}\mathrm{EMA}_{t}\left(\eta\right)\\[10.0pt] or\\[10.0pt] \hat{\mathrm{ARP}}_{t+1}=\left(1-\rho\right)\hat{\mathrm{ARP}}_{t}+\rho\Sigma^{-1}_{t}C^{-0.5}\mathrm{MACD}_{t}\left(\eta_{1},\eta_{2},\eta_{3},\omega_{1},\omega_{2},\omega_{3}\right)\\ \end{array}\right.$ (12)
${\mathrm{ARP}}_{t+1}=\frac{\hat{\mathrm{ARP}}_{t+1}}{\sqrt{\hat{\mathrm{ARP}}_{t+1}^{\prime}{\Sigma_{t}}C{\Sigma_{t}}\hat{\mathrm{ARP}}_{t+1}}}$ (13)
3.4 Data and different simulated parameters
The simulation starts on the 25th of may 1990 and stops on the 7th of december 2023. We used daily returns from 70 futures instruments in stock indices, bonds, fx and commodities futures. The description is in the appendix A.
We tested the different indicators applying the $\mathrm{ARP}$ formula Eq.13. The different parameters are described in Tab 3 in the appendix B.
Fig.4 displays the sensitivities of these indicators to past daily returns. $\mathrm{MACD}$ enables to put more weights on very older returns as we expected.
Figure 4: Different sensitivities to past daily returns for indicators ($\mathrm{MACD}$, $\mathrm{EMA}$) all based on 20, 80 and 400 days time scales. we have $\mathrm{ARP}\left(0\times 20,80,0.2\right)$, $\mathrm{ARP}\left(0\times 20,80,0.4\times 400\right)$, $\mathrm{ARP}\left(20,80,0.2\times 400\right)$, $\mathrm{ARP}\left(20,80,0.4\times 400\right)$,$\mathrm{ARP}\left(80\right)$, $\mathrm{ARP}\left(400\right)$
4 Empirical results
Here we first interpret the simulated empirical Sharpe ratio when applying the $\mathrm{ARP}$ portfolio with the diffent indicators. We find interesting results which challenge traditional recipes.
4.1 Grebenkov’s model empirical validation
Interestingly, the empirical simulation fits pretty well the theorical formula of the Sharpe ratio Eq.4 derived in Grebenkov and Serror (2014) with the following parameters $\lambda=1/180$ and $\beta_{0}=0.12$. The Fig.5 is very impressive and the empricial fits should validate the Gebenkov’s model to describe trends. Moreover the fit appears to be more robust than the variogram (Fig.2) of the Dow Jones Index. That is the main result of the paper.
As a first consequence the parameter of $112$ days ($\eta_{\text{opt}}=\frac{1}{112}$) for simple $\mathrm{EMA}$ is the optimal parameter to get the optimal Sharpe ratio when not accounting cost of trading which are very small at that trading frequency. Lamperiere et al (2014) found a different result, with a faster optimal parameter at 50 days for the EMA, but very similar Sharpe ratios across EMAs ranging from 1 to 5 months, showing an almost flat curve with no clear optimum. The difference with our results may be explained by the choice made in Lamperiere et al (2014) to determine positions based not on the linear magnitude of the EMA signals, but solely on their sign. Additionally, they did not use the ARP portfolio construction, which involves inverting the square root of the correlation matrix—a method that explains a significant part of the Sharpe ratio, as shown in Benichou at al (2017). For these two reasons, their Sharpe ratio for the post-2000 period was measured at 0.85, lower than our 1.2, and their optimal EMA signal corresponded to a faster timescale than ours.
As a second consequence of the very good fit with the theoritical formula, modeling trend through mean-reversion process using only one relaxation time and not an multi-time scales one appears to be a good solution as the fit is more than correct. That is particulary unexpected as market is known to have a multi time scales property: For example the relaxation of volatility is known to have muti-time scales, investors are expected to have different horizons of time and different horizons of analysis. As a consequence we can wonder wether the usual recipe to take into account of a multitude of different indicators is justified to claim having the most robust and the most refined signal.
We can also note that $\beta_{0}=0.12$ is slightly higher than $0.08$, the parameter measured for the Dow Jones over the past 100 years by Grebenkov and Serror (2014). This difference may seem minor, but since the Sharpe ratio depends on $\beta_{0}^{2}$, it results in a Sharpe ratio that is 2.25 times higher when applying the strategy to a universe of 70 underlying assets instead of the Dow Jones (assuming the trend strength over the past 30 years was similar to that of the past 100 years). This further confirms the importance of measuring the implied autocorrelation parameter based on a strategy invested in a large universe, as it leads to more accurate and agregated estimates.
We can also note that $\lambda$ is estimated to $\frac{1}{180}$ instead of $0.011$ in the case of the Dow Jones. Our analysis is that market behavior should have changed in the last 100 years and we believe that our fit appears more robust than a simple variogram.
We can see based on the Tab 4 in appendix C that empirical Sharpe is $1.24$ for $\mathrm{ARP}\left(120\right)$ with one time scale and $1.18$ for $\mathrm{MACD}\left(20,120,0.4\times 400\right)$ with 3 time scales. $\mathrm{MACD}$ could not be justified as additional time scales does not bring significant improvement. Also Sharpe ratio is not so sensitive to the parameters arround the optimal as expected: Sharpe is $1.25$ for $\mathrm{ARP}\left(100\right)$ and $1.21$ for $\mathrm{ARP}\left(150\right)$.
Figure 5: Empirical Sharpe ratio based on the whole period 1990-2023 for trend following strategies using $\mathrm{ARP}$ (Eq.13 ) applied to $\mathrm{EMA}$ as signal v.s. $\eta$ the parameter of the $\mathrm{EMA}$ (Eq.6) and the theoretical equation Eq.4. The theoritical model is fitted with parameters $\beta_{0}=0.12$ and $\lambda=1/180$ so that $\eta_{\text{opt}}=\frac{1}{112}$ Empirical measurements are displayed in the tab 4.
5 Replication of a simple $\mathrm{EMA}$ by a mixture of bollinger bands $\mathrm{BB}$
The Bollinger Bands indicator, $\mathrm{BB}$, applies a double heavy-side function to an $\mathrm{SMA}$ with a width $\delta$. This indicator is nonlinear and therefore complex, with sensitivities to past returns that depend on the price path. However, the usual approach is to use a signal composed of a mixture of many Bollinger Bands indicators with different parameters, making the signal less path-dependent and increasingly robust. Here, we aim to demonstrate that a simple $\mathrm{EMA}$ signal can be replicated through a complex mixture of a large number of $\mathrm{SMA}$, which can, in turn, be decomposed into a large number of $\mathrm{BB}$ Bollinger Bands (Eq.16). This explains why it is common, as shown in Fig.6 when replication the optimal $\mathrm{EMA}$ with 112 days, to display indicator weights following a bell-shaped curve centered around 200 days, while explaining to investors that the signal contains both short-term and long-term indicators.
$\mathrm{EMA}_{t}\left(\eta\right)=\frac{r_{t}+\left(1-\eta\right)r_{t-1}+\left(1-\eta\right)^{2}r_{t-2}+....+\left(1-\eta\right)^{n}r_{t-n}+....}{1+\left(1-\eta\right)+\left(1-\eta\right)^{2}+....+\left(1-\eta\right)^{n}+...}$ (14)
$\mathrm{EMA}_{t}\left(\eta\right)=\frac{...+\left[\left(1-\eta\right)^{n-1}\eta\right]\left(n-1\right)\mathrm{SMA}_{t}\left(n-1\right)+...+\left[\left(1-\eta\right)\eta\right]\mathrm{SMA}_{t}\left(1\right)}{1+\left(1-\eta\right)+\left(1-\eta\right)^{2}+....+\left(1-\eta\right)^{n}+...}$ (15)
[…]
Figure 6: weigths of a mixture of $\mathrm{BB}$ Vs size of the window to replicate the $\mathrm{EMA}\left(\frac{1}{112}\right)$ which is derived from Eq.16 corrected to pick different parameters of $\mathrm{BB}$ uniformely into lag in logarithm
5.1 Correlations between indicators
Fig.7 in the Appendix C shows that the indicators with different parameters give very correlated strategies which confims it makes not so appealing the taditional approach of using a basket of many indicators as signal. $\mathrm{ARP}\left(80\right)$ is correlated to $\mathrm{ARP}\left(150\right)$ with a coefficient of $0.96$. That is nevertheless pretty interesting to see so strong correlations which could be explained by a common factor somewhere, which could be the object of an additional research. Morevever the $\mathrm{ARP}\left(120\right)$ based on a simple $\mathrm{EMA}$ is very close with correlation from 1 to 0.99 to the new indicator $\mathrm{MACD}\left(20,120,0\times 400\right)$ or $\mathrm{MACD}\left(20,120,0.4\times 400\right)$ we introduced based on 3 different times scales to increase weights for very old past returns and decrease weights on very recent ones. So it makes this rafinement not justified again.
5.2 Conclusion
Grebenkov’s model for describing trends is empirically validated, as its pretty complicated theoretical formula for determining the Sharpe ratio based on the $\mathrm{EMA}$ parameter fits impressively well with empirical data. The best fit is obtained using the theoretical model parameters $\lambda=1/180$ and $\beta_{0}=0.12$. As a consequence the parameter of $112$ days for simple $\mathrm{EMA}$ is the optimal parameter to get the optimal Sharpe ratio. It is quite surprising that a single $\mathrm{EMA}$ is optimal for capturing trends, as one would expect different time scales for different types of investors. However, there are likely much shorter time scales, on the order of a few days, but they have no significant impact on a medium-frequency strategy. The conclusion is that using a complex mixture of sophisticated indicators is unnecessary when the $\mathrm{EMA}$ alone provides a perfect fit—proving that simplicity can indeed be beautiful.
Appendix A Data
Commodities
Brent Crude (IFEU \$/bbl)
Cocoa (IFUS \$/mt)
Coffee (IFUS \$/lbs)
Corn (CBT \$/bu)
Cotton #2 (IFUS \$/lbs)
Crude Oil WTI (NYM \$/bbl)
ECX EUA (IFEU EUR/t)
Feeder Cattle (CME \$/lbs)
Gasoil (IFEU \$/mt)
Gold (NYM \$/ozt)
Hard Red Wtr Wheat (CBT \$/bu)
High Grade Copper (NYM \$/lbs)
Iron Ore 62% Fe, CFR China (TSI) (NYM \$/mt)
Lean Hogs (CME \$/lbs)
Live Cattle (CME \$/lbs)
Lumber (CME \$/bft)
Milling Wheat (LIF EUR/t)
Natural Gas (NYM \$/mmbtu)
NY Harb RBOB (NYM \$/gal)
NY Harbor ULSD (NYM \$/gal)
Oats (CBT \$/bu)
Orange Juice (IFUS \$/lbs)
Palladium (NYM \$/ozt)
Platinum (NYM \$/ozt)
Rough Rice (CBT \$/cwt)
Rubber RSS3 (TKT JPY/kg)
Silver (NYM \$/ozt)
Soybean Meal (CBT \$/t)
Soybean Oil (CBT \$/lbs)
Soybeans (CBT \$/bu)
Sugar #11 (IFUS \$/lbs)
Wheat (Chicago) - Contract
Table 1: List of the 32 commodities futures traded in the USA or in Europe.
Stock indices Bond indices FOREX
AMSTERDAM EOE Idx
S&P Midcap 400 Idx e-mini
Russell 2000 Idx e-mini
Cac 40
Dax
Ftse 100
STOXX Europe 600 Index Futures
Hang Sen
Mib S&p-mif
Nikkei 225 Osaka
Topix
Kospi 200
Ibex 35
Mini MSCI Emerging Markets Index Future
Nasdaq E-mini
S&P 500 e-mini
Dj Euro Stoxx
S&P Canada 60-ME
SPI 200 Idx
Mini Dow Futures
BUND 10Yr
CAD Bond 10Yr
Bobl
Schatz
Long-term Euro-btp
Euro-buxl Futures
LONG Gilt 10Yr
10yr Fr Gov Bond
US T-NOTE 5Yr
JGB 10Yr
US T-NOTE 10Yr
US T-Note 2Yr
Ultra T-Bonds Combined
AUD/USD Fut.
GBP/USD Fut.
CAD/USD Fut.
EUR/USD Fut.
JPY/USD Fut.
MXN/USD Fut.
NZD/USD Fut.
CHF/USD Fut.
Table 2: List of the 41 instruments among currenies, equity indices and Bonds.
Appendix B Parameters
portfolio comment indicator type
$\mathrm{ARP}\left(20\right)$ for 20 days as relaxation time for the $\mathrm{EMA}$ when applying $\mathrm{EMA}\left(\frac{1}{20}\right)$
$\mathrm{ARP}\left(50\right)$ when applying $\mathrm{EMA}\left(\frac{1}{50}\right)$
$\mathrm{ARP}\left(80\right)$ when applying $\mathrm{EMA}\left(\frac{1}{80}\right)$
$\mathrm{ARP}\left(100\right)$ when applying $\mathrm{EMA}\left(\frac{1}{100}\right)$
$\mathrm{ARP}\left(120\right)$ when applying $\mathrm{EMA}\left(\frac{1}{120}\right)$
$\mathrm{ARP}\left(150\right)$ when applying $\mathrm{EMA}\left(\frac{1}{150}\right)$
$\mathrm{ARP}\left(180\right)$ when applying $\mathrm{EMA}\left(\frac{1}{180}\right)$
$\mathrm{ARP}\left(400\right)$ when applying $\mathrm{EMA}\left(\frac{1}{400}\right)$
$\mathrm{ARP}\left(1000\right)$ when applying $\mathrm{EMA}\left(\frac{1}{1000}\right)$
$\mathrm{ARP}\left(20,120,0\times 400\right)$ when applying $\mathrm{MACD}\left(\frac{1}{20},\frac{1}{120},\frac{1}{400},\omega_{1},1,0\right)$
$\mathrm{ARP}\left(0\times 20,120,0\times 400\right)$ when applying $\mathrm{MACD}\left(\frac{1}{20},\frac{1}{120},\frac{1}{400},0,1,0\right)$
$\mathrm{ARP}\left(20,120,0.2\times 400\right)$ when applyingr $\mathrm{MACD}\left(\frac{1}{20},\frac{1}{120},\frac{1}{400},\omega_{1},1,0.2\right)$
$\mathrm{ARP}\left(20,120,0.4\times 400\right)$ when applying $\mathrm{MACD}\left(\frac{1}{20},\frac{1}{120},\frac{1}{400},\omega_{1},1,0.4\right)$
$\mathrm{ARP}\left(20,90,0.3\times 400\right)$ when applying $\mathrm{MACD}\left(\frac{1}{20},\frac{1}{90},\frac{1}{400},\omega_{1},1,0.3\right)$
$\mathrm{ARP}\left(20,80,0.3\times 400\right)$ when applying $\mathrm{MACD}\left(\frac{1}{20},\frac{1}{80},\frac{1}{400},\omega_{1},1,0.3\right)$
$\mathrm{ARP}\left(20,80,0.2\times 400\right)$ when applying $\mathrm{MACD}\left(\frac{1}{20},\frac{1}{80},\frac{1}{400},\omega_{1},1,0.2\right)$
$\mathrm{ARP}\left(20,80,0.4\times 400\right)$ when applying $\mathrm{MACD}\left(\frac{1}{20},\frac{1}{80},\frac{1}{400},\omega_{1},1,0.4\right)$
Table 3: Different parameters fo the trend indicator
Appendix C Results
portfolio indicator type Gross Sharpe ratio average holding period (days))
$\mathrm{ARP}\left(20\right)$ $\mathrm{EMA}$ 1.079535 38
$\mathrm{ARP}\left(50\right)$ $\mathrm{EMA}$ 1.189320 60
$\mathrm{ARP}\left(80\right)$ $\mathrm{EMA}$ 1.240349 74
$\mathrm{ARP}\left(100\right)$ $\mathrm{EMA}$ 1.244945 81
$\mathrm{ARP}\left(120\right)$ $\mathrm{EMA}$ 1.235455 88
$\mathrm{ARP}\left(150\right)$ $\mathrm{EMA}$ 1.207496 96
$\mathrm{ARP}\left(180\right)$ $\mathrm{EMA}$ 1.172569
$\mathrm{ARP}\left(400\right)$ $\mathrm{EMA}$ 0.955223 132
$\mathrm{ARP}\left(1000\right)$ $\mathrm{EMA}$ 0.633678 155
$\mathrm{ARP}\left(0\times 20,120,0\times 400\right)$ $\mathrm{MACD}$ 1.235455 88
$\mathrm{ARP}\left(20,120,0.2\times 400\right)$ $\mathrm{MACD}$ 1.203418 97
$\mathrm{ARP}\left(20,120,0.4\times 400\right)$ $\mathrm{MACD}$ 1.176466
$\mathrm{ARP}\left(20,90,0.3\times 400\right)$ $\mathrm{MACD}$ 1.214172 89
$\mathrm{ARP}\left(20,80,0.3\times 400\right)$ $\mathrm{MACD}$ 1.218031 85
$\mathrm{ARP}\left(20,80,0.2\times 400\right)$ $\mathrm{MACD}$ 1.228186 82
$\mathrm{ARP}\left(20,80,0.4\times 400\right)$ $\mathrm{MACD}$ 1.206864 87
Table 4: Gross sharpe ratio based on the whole period 1990-2023
Figure 7: Empirical correlation on the whole period 1990-2023 between ARP models with different trend indicators parameters

### v2

@@
 Abstract
-Our empirical results, illustrated in Fig.5, show an impressive fit with the pretty complex theoritical Sharpe formula of a Trend following strategy depending on the parameter of the signal, which was derived by Grebenkov and Serror (2014).
+Our empirical results show an impressive fit with the pretty complex theoritical Sharpe formula of a Trend following strategy depending on the parameter of the signal, which was derived by Grebenkov and Serror (2014).
 That empirical fit convinces us that a mean-reversion process with only one time scale is enough to model, in a pretty precise way, the reality of the trend-following mechanism at the average scale of CTAs and as a consequence, using only one simple $\mathrm{EMA}$, appears optimal to capture the trend.
 As a consequence, using a complex basket of different complex indicators as signal, do not seem to be so rational or optimal and exposes to the risk of cherry-picking.
-April 15, 2025
+May 17, 2025
 I thank Prof Campbell Harvey, Robert Kosowsky, Valeriy Zakamulin, Denis Grebenkov and Raphael Douady for their feedbacks and valuable insights.
 1 Introduction
+Practitioners of the trend-following strategies usually optimize, when replaying the historical scenarii through past returns, their strategies empirically by selecting parameters that maximize the simulated reward-to-risk ratio, typically measured by the Sharpe ratio.
+However, this empirical approach can lead to overfitting, especially when optimization is performed on the same historical dataset used for evaluation.
+An alternative is to model the returns of the underlying assets using stochastic processes.
+In some cases, this allows the Sharpe ratio to be derived mathematicaly through a formula depending on the parameters of the strategy.
+When this is possible, we refer to it as the theoretical Sharpe ratio, in contrast to the empirical one.
+It is also possible to determine the set of parameters that maximizes the theoretical Sharpe ratio—this is known as the theoretical optimal Sharpe ratio, as opposed to the common empirical optimum obtained through backtesting.
+A theoritical Sharpe ratio is always interesting if the model deriving the returns is realistic enough, as it allows to validate and optimize a strategy without the risk of overfiting.
+The main objective of this paper is to empirically validate the theoretical Sharpe ratio model proposed by Grebenkov and Serror (2014) and therefore their assumption to model trends as gaussian mean reversion stochastic processes with only one time scale.
+2 Literature
 The usual recipe to determine a portfolio in the CTA’s industry is to use a blend of many different technical indicators as signal for each underlying instrument (most often individual indicator gives either a long or a short position).
@@
 He also employs a two-state regime-switching model (bull and bear regimes), a widely accepted framework for stock returns (Timmermann (2000); Fruhwirth-Schnatter (2006); Giner and Zakamulin (2023)), using a semi-Markov model to detect negative autocorrelations over very long time scales.
-He argues that, within this framework, the MACD indicator with three different time scales could be adapted to construct the optimal portfolio.
-However, mean reversion at larger scales is unconvincing, and the two-state regime does not seem natural for modeling trend-following mechanisms, which are better explained by herding behavior rather than by an external guiding force.
+He argues that, within this framework, the MACD indicator with three different time scales could be adapted to construct the optimal portfolio but obtaining statistically significant results remains difficult.
+So mean reversion at larger scales remains unconvincing.
+However the two-state regime does not seem natural for modeling trend-following mechanisms, which are better explained by herding behavior rather than by an external guiding force.
 Models in which agents partially imitate each other are more likely to produce continuous regime changes rather than abrupt shifts between two states.
@@
 3 Theory from Grebenkov and Serror (2014)
-Grebenkov and Serror (2014), who solved the correct optimization problem—maximizing the Sharpe ratio at the portfolio level while ensuring that positions remain linearly dependent on signals—introduced a diffusive process with a positive autocorrelation model for returns $r_{i,t}$ of the $i$th instrument among the N ones to describe trends in financial market prices , as formulated in Eq.1.
+To derive the theoretical Sharpe ratio for a trend-following strategy, we need a model that describes returns and trends.
+A classical one-dimensional model used to capture trends in an underlying asset assumes that returns consist of two components: a random, independent and identically distributed (i.i.d.) component, and a small bias term.
+This bias can be modeled using an auto regressive process of order 1, AR(1), which represents mean reversion process around zero with only one time scale.
+The AR(1) process is characterized by one time scale parameter—indicating the average time needed to revert to zero—and a standard deviation that determines the level of noise.
+AR(1) is very similar if time lag is infinitesimally small to the Orstein-Uhlenbeck process which is a mean reversion usual and natural process in complex systems with only one time scale.
+This classical one-dimensional model can be easily extended to multiple dimensions.
+In the multivariate case, two covariance matrices are required to account for correlations.
+The first matrix captures the correlations between the random components of the different underlyings, while the second matrix reflects the correlations among the AR(1) processes associated with each underlying.
+The model is formally described in Eq.
+1, using the same notation as in Grebenkov and Serror (2014).
+In that equation the returns $r_{i,t}$ of the $i$th instrument among the N is described in two components: The first $\epsilon_{i,t}$ is a noise which is not autocorrelated while the second $\beta\sum_{k=1}^{t-1}{\left(1-\lambda\right)^{t-1-k}\xi_{i,k}}$ is an AR(1) process which is autocorrelated where $\xi_{i,k}$ is a noise which is not autocorrelated.
+Ultimately Eq.1 describes the returns of a diffusive process with a positive autocorrelation model for returns $r_{i,t}$: The weights $\left(1-\lambda\right)^{t-1-k}$ are declining exponentialy.
+The parameter $\lambda$ describes the inverse of the relaxation time of the mean reversion AR(1) process describing the short bias, and $\beta$ represents the strength of that short stochastic bias in the returns.
+The noises $\xi_{i,k}$ are independent in the $k$ axis but may be correlated with $i$ axis (Valeyre (2024) introduces specific forms of that matrix, leading to particular optimal portfolios).
+The noises $\epsilon_{i,t}$ are independent in the $t$ axis but may be correlated with $i$ axis with a very close correlation matrix between returns.
+Through that model, the trend is $\beta\sum_{k=1}^{t-1}{\left(1-\lambda\right)^{t-1-k}\xi_{i,k}}$ and it follows an auto regressive model of order 1, AR(1).
 $r_{i,t}=\epsilon_{i,t}+\beta\sum_{k=1}^{t-1}{\left(1-\lambda\right)^{t-1-k}\xi_{i,k}}$ (1)
-The weights $\left(1-\lambda\right)^{t-1-k}$ are declining exponentialy.
-The parameter $\lambda$ describes the inverse of the relaxation time of the autocorrelation, and $\beta$ represents the strength of that autocorrelation component in the returns.
-The noises $\xi_{i,k}$ are independent in the $k$ axis but may be correlated with $i$ axis (Valeyre (2024) introduces particular cases).
-The noises $\epsilon_{i,t}$ are independent in the $t$ axis.
-Through that model, the trend is $\beta\sum_{k=1}^{t-1}{\left(1-\lambda\right)^{t-1-k}\xi_{i,k}}$ follows an auto regressive model of order 1 or an Orstein-Uhlenbeck process if time lag is infinitesimally small which is a mean reversion usual and natural process in complex systems with only one time scale.
 Through variogram measurements on the Dow Jones index, they estimated that $\lambda=0.01$ and $\beta_{0}=0.1$ (Fig.2).
-The fit appears suspiciously perfect, considering the expectation of multiple time scales with long memory among investors.
+The fit appears suspiciously perfect, considering the expectation of multiple time scales with long memory among investors as well as the potential presence of mean-reversion patterns at longer horizons that are not captured.
 Additionally, the result is not entirely convincing, as the 100-year period may be heterogeneous, and the fit is less accurate at shorter time scales, where measurements should, in principle, be less noisy.
@@
 Figure 2: Capture of Grebenkov and Serror (2014). Variogram applied to Dow Jones (1900-2012)to fit $\beta_{0}=0.08$ and $\lambda=0.011$ parameters
+Grebenkov and Serror (2014) solved the correct theoritical optimization problem—maximizing the Sharpe ratio at the portfolio level while ensuring that positions remain linearly dependent on signals when assuming that returns are well described by Eq.1.
+They considered only the one-dimensional case, but later, Grebenkov and Serror (2015) and Valeyre (2024) extended the analysis to the multi-dimensional setting.
 Instead of focusing on the best predictor or on the optimal portfolio with suboptimal constraints—such as the possibility of being either long with the same risk or flat, which is unfortunately common in the literature—Grebenkov and Serror (2014) assumes that positions should better depend linearly on the signals, which is a natural approach which is in agreement with the Markowitz solution.
@@
 Thus, Grebenkov and Serror (2014) introduced substantial novelty, even though it is not yet widely considered by most academics specializing in trend-following.
-In detail, they derived an explicit formula in Eq.3 for the optimal parameter ($\eta_{\text{opt}}$) of the Exponential Moving Average ($\mathrm{EMA}$) applied to returns as a trend indicator for a trend-following strategy (Eq.6 mathematically defines later this indicator).
-Additionally, they derived in Eq.4 the theoretical Sharpe ratio function ($\mathrm{SHARPE}\left(\eta\right)$), assuming returns are generated through their autocorrelation model (1), as a function of $\eta$ and $\lambda$, while also considering the inclusion of trading costs, denoted as $\theta$.
-$\beta=\beta_{0}\sqrt{\lambda\left(2-\lambda\right)}$ (2)
-$\eta_{\text{opt}}=\lambda\sqrt{1+2\frac{\beta^{2}_{0}}{\lambda}}$ (3)
-$\mathrm{SHARPE}\left(\eta\right)=\frac{\beta^{2}_{0}\sqrt{2\eta}-\frac{2}{\pi}\theta\sqrt{\eta}\left(\lambda+\eta\right)}{\sqrt{\left(\lambda+\eta\right)^{2}+2\beta^{2}_{0}\left(\lambda+\eta\right)}}$ (4)
+In detail, they derived an explicit formula, in Eq.
+3, the theoretical Sharpe ratio (${\mathbf{S}}\left(\eta\right)$), when assuming returns are generated through their autocorrelation model in Eq.
+1, as a function of $\eta$ (the smoothing parameter of the trend indicator $EMA$ which is defined mathematically later in Eq.6 ) , $\lambda$ (the inverse of the time scale of the mean reversion process that describes the trends in Eq.
+1) and $\beta$ (the weight of the trend component in the returns Eq.
+1) while also considering the inclusion of trading costs, denoted as $\theta$.
+Equation Eq.3 is in fact a remarkably elegant formula.
+As expected, the theoretical Sharpe ratio depends on $\beta$, the weight of the trend in the returns, in an approximately proportional manner when $\beta_{0}^{2}$ is significantly larger than $\frac{1}{\lambda}$ and when $\eta$ remains close to $\lambda$.
+In that case, the Sharpe ratio from Eq.3 can be approximated by $\beta_{0}\sqrt{\frac{1}{2}}$ or $\beta_{0}\sqrt{\frac{255}{2}}$ in the annualized version of the Sharpe.
+Fig.3 illustrates the very complex theoritical formula Eq.3 through a simple graph which enable to undertsand its sensitivity against the parameter of the $\mathrm{EMA}$.
+In reality, Eq.
+3 was derived for the one-dimensional case only.
+Nevertheless, we can reasonably expect the equation to hold in the multi-dimensional setting.
+In that case, Eq.
+2 must be adjusted to redefine $\beta_{0}$.
+To be consistent with the observed increase of the Sharpe ratio with the size of the universe $N$, we can, under strong approximations, expect $\beta_{0}$ to be related to $\frac{\beta}{\sqrt{\lambda\left(2-\lambda\right)}}\sqrt{\frac{N}{1+\left(N-1\right)\rho^{2}}}$ where $\rho^{2}$ is the average squared correlation in the universe.
+This expression extends Eq.
+2 using Eq.
+(40) from Grebenkov and Serror (2015).
+Additionally they derived in Eq.4, the optimal smoothing parameter ($\eta_{\text{opt}}$) of the Exponential Moving Average ($\mathrm{EMA}$) applied to returns as a trend indicator for a trend-following strategy which is supposed to yield to the optimal Sharpe ratio.
 It is interesting to observe from Eq.4 that $\frac{1}{\eta_{\text{opt}}}$, the time scale of the signal $\mathrm{EMA}$ generating the optimal strategy is always shorter than $\frac{1}{\lambda}$, the time scale of the best predictor $\mathrm{EMA}$ (or the optimal indicator in the sub optimal optimization problem of Zakamuli and Giner (2024)) and that of the autocorrelated process of returns.
@@
 As a conclusion, theoretically the parameters of the signals should be adjusted when the universe is increased.
-Fig.3 illustrates the very complex theoritical formula Eq.4 through a simple graph which enable to undertsand its sensitivity against the parameter of the $\mathrm{EMA}$.
-[…]
+$\beta=\beta_{0}\sqrt{\lambda\left(2-\lambda\right)}$ (2)
+${\mathbf{S}}\left(\eta\right)=\frac{\beta^{2}_{0}\sqrt{2\eta}-\frac{2}{\pi}\theta\sqrt{\eta}\left(\lambda+\eta\right)}{\sqrt{\left(\lambda+\eta\right)^{2}+2\beta^{2}_{0}\left(\lambda+\eta\right)}}$ (3)
+$\eta_{\text{opt}}=\lambda\sqrt{1+2\frac{\beta^{2}_{0}}{\lambda}}$ (4)
+Figure 3: Capture of Grebenkov and Serror (2014).
+Theoritical Sharpe (${\mathbf{S}}$ from Eq.
+3) of the trend following strategy depending on $\eta$ the parameter of the $\mathrm{EMA}$.
+$\beta_{0}=0.1$ and $\lambda=0.01$ for different cost of trading $\theta$
 4 Description of the empirical analysis
-3.1 Single $\mathrm{EMA}$ indicator
-The $\mathrm{EMA}$ at the time $t$ is a vector of $N$ exponential moving averages of normalized returns, one for each of the $N$ instruments.
-The $EMA$ of the instrument $i$ at time $t+1$ is defined by Eq.6.
+To validate Grebenkov’s theory (Eq.
+3), introduced in Sec 3 and, therefore, the diffusive process in Eq.
+1 which is assumed for the underlying assets, we conducted empirical backtests of the Agnostic Risk Parity (ARP) strategy introduced by Benichou at al (2017).
+This strategy constructs a rotationally invariant portfolio that applies trend-following while allocating equal unconditional risk to each eigenvector of the covariance matrix.
+While rotational invariance is a desirable theoretical property, the equal unconditional risk allocation among eigenvectors is a more desired property for asset manager.
+Moreover, Valeyre (2024) demonstrated that the ARP portfolio is optimal—achieving the highest Sharpe ratio—under the assumption that the covariance of the $\xi_{i,k}$ terms in Eq.
+1 follows a particular random matrix structure, which is considered realistic.
+Implementing the ARP portfolio is relatively straightforward and consists of normalizing the trend indicators by the inverse square root of the correlation matrix, rather than by the full inverse as in the classical Markowitz framework.
+We decided to focus primarily on exponential moving averages ($\mathrm{EMA}$) as trend indicators to feed the Agnostic Risk Parity (ARP) strategy, as they are theoretically optimal according to Grebenkov and Serror (2014); Zakamuli and Giner (2024), assuming the validity of Eq.
+1.
+Accordingly, we tested various $\mathrm{EMA}$ configurations and measured how the empirical Sharpe ratio of the resulting portfolio depends on the EMA parameter, in line with Grebenkov’s theoretical predictions.
+Additionally, we explored more complex trend indicators based on multiple time scales—such as the Moving Average Convergence Divergence ($\mathrm{MACD}$)—to empirically validate whether the $\mathrm{EMA}$ remains the optimal choice.
+We backtested the Agnostic Risk Parity (ARP) strategy on a global universe including Futures on commodities, fx, stock indices and bonds on a period 25th of may 1990 and stops on the 7th of december 2023.
+4.1 Renormalized exponential moving average indicator ($\mathrm{EMA}$)
+We used only a normalized version of the $EMA$ to account for the dependence of volatility on both time and instrument.
+This normalization allows for meaningful comparisons of trend indicators across instruments, time periods, and even different smoothing parameters.
+At time $t$, the $\mathrm{EMA}$ is a vector of $N$ exponential moving averages of normalized returns, one for each of the $N$ instruments.
+The $EMA$ of instrument $i$ at time $t+1$ is defined by $\phi$ in Eq.
+6, where $r_{i,t+1}$ denotes the return of instrument $i$ at time $t+1$.
+The parameter $\sigma$ represents the standard deviation of returns used to normalize them, and $\eta$ is the smoothing parameter of the $EMA$ as well as of the return normalization.
+$\sigma^{2}$ can be interpreted as an exponential moving average of squared returns, and the indicator $\phi$ corresponds to an exponential moving average of returns scaled by their volatility..
 The incrementation is on daily basis but it could be generalized to minutes returns.
@@
 […]
-The indicator $\mathrm{EMA}\left(150\right)$ is defined by the $\mathrm{EMA}_{i,t+1}\left(\eta\right)$ when $\eta=\frac{1}{150}$ is applied to every underlying $i$ and time $t$.
-We use $\sqrt{\eta}$ in Eq.6 so that the std of $\mathrm{EMA}_{i,t+1}\left(\eta\right)$ is 1 theoritically if we have a random walk.
-3.2 $\mathrm{MACD}$ as a 3 time scales $\mathrm{EMA}$ indicator
-Inspired by $\mathrm{MACD}$, we introduced a combination of $\mathrm{EMA}$ applying the Eq.7 while determining $\omega_{1}$ so that the derivative of the sensitivity to past daily returns is at 0 at the lag 0 and replicate curves in Fig.4
-$\mathrm{MACD}_{i,t}\left(\eta_{1},\eta_{2},\eta_{3},\omega_{1},\omega_{2},\omega_{3}\right)=\omega_{1}\mathrm{EMA}_{i,t}\left(\eta_{1}\right)+\omega_{2}\mathrm{EMA}_{i,t}\left(\eta_{2}\right)+\omega_{3}\mathrm{EMA}_{i,t}\left(\eta_{3}\right)$ (7)
+The indicator $\mathrm{EMA}\left(150\right)$ is defined by the $\phi_{i,t+1}\left(\eta\right)$ when $\eta=\frac{1}{150}$ is applied to every underlying $i$ and time $t$.
+We use $\sqrt{\eta}$ in Eq.6 so that the std of $\phi_{i,t+1}\left(\eta\right)$ is 1 theoritically if $r$ are well approximated by unautocorrelated returns so that $\phi$ is a normalized $EMA$ indicator.
+4.2 Moving Average Convergence Divergence indicator ($\mathrm{MACD}$) as a 3 time scales $\mathrm{EMA}$ indicator
+Inspired by $\mathrm{MACD}$, we introduced a combination of $\mathrm{EMA}$ applying the Eq.7 while determining $\omega_{1}$ so that the derivative of the sensitivity to past daily returns is at 0 at the lag 0 and replicate curves in Fig.4.
+The idea behind this indicator was to reduce the weight of returns at very short time scales—where mean reversion is possible—and to increase the weight of returns at longer time scales, in order to test a multi-timescale memory effect for trend-followers.
+$\mathrm{MACD}_{i,t}\left(\eta_{1},\eta_{2},\eta_{3},\omega_{1},\omega_{2},\omega_{3}\right)=\omega_{1}\phi_{i,t}\left(\eta_{1}\right)+\omega_{2}\phi_{i,t}\left(\eta_{2}\right)+\omega_{3}\phi_{i,t}\left(\eta_{3}\right)$ (7)
 The derivative at zeros yields to Eq.8.
@@
 The vector $\Sigma$ consists of N values representing the standard deviations estimated using a 40-day exponential moving average applied to daily returns.
-Next we expressed the positions vector as a linear function of the signals ($\mathrm{EMA}$ or $\mathrm{MACD}$) applying a normalisation that involves the inverse of the square root of the estimated correlation matrix and volatilities .
-This follows the formulation of the agnostic risk parity portfolio (${\mathrm{ARP}}$) introduced by Benichou at al (2017)
-We then used the parameter $\rho=\frac{1}{20}$ for portfolio smoothing as specified in Eq.LABEL:eq:arp2.
-Finally, wer applied a resizing process in Eq.13 to target a constant volatility for the final positions ${\mathrm{ARP}}$ which is a vector of N weights at time $t+1$.
-$\left\{\begin{array}[]{l}\hat{\mathrm{ARP}}_{t+1}=\left(1-\rho\right)\hat{\mathrm{ARP}}_{t}+\rho\Sigma^{-1}_{t}C^{-0.5}\mathrm{EMA}_{t}\left(\eta\right)\\[10.0pt] or\\[10.0pt] \hat{\mathrm{ARP}}_{t+1}=\left(1-\rho\right)\hat{\mathrm{ARP}}_{t}+\rho\Sigma^{-1}_{t}C^{-0.5}\mathrm{MACD}_{t}\left(\eta_{1},\eta_{2},\eta_{3},\omega_{1},\omega_{2},\omega_{3}\right)\\ \end{array}\right.$ (12)
+Next we expressed the positions vector as a linear function of the signals ($\mathrm{EMA}$ through $\phi$ or $\mathrm{MACD}$) applying a normalisation that involves the inverse of the square root of the estimated correlation matrix and volatilities .
+This follows the formulation of the agnostic risk parity portfolio (${\mathrm{ARP}}$) introduced by Benichou at al (2017).
+We then used the parameter $\rho=\frac{1}{20}$ for portfolio smoothing as specified in Eq.LABEL:eq:arp2 which is an very easy solution to reduce trading cost (usually, trend indicators include rules to reduce trading costs.
+While the $EMA$ does not explicitly aim to avoid excessive trading, smoothing the portfolio has proven to be both simple and effective.
+Benichou at al (2017) recommended optimization under constraints to better manage trading costs).
+Finally, we applied a resizing process in Eq.13 to target a constant volatility for the final positions ${\mathrm{ARP}}$ which is a vector of N weights at time $t+1$.
+$\left\{\begin{array}[]{l}\hat{\mathrm{ARP}}_{t+1}=\left(1-\rho\right)\hat{\mathrm{ARP}}_{t}+\rho\Sigma^{-1}_{t}C^{-0.5}\phi_{t}\left(\eta\right)\\[10.0pt] or\\[10.0pt] \hat{\mathrm{ARP}}_{t+1}=\left(1-\rho\right)\hat{\mathrm{ARP}}_{t}+\rho\Sigma^{-1}_{t}C^{-0.5}\mathrm{MACD}_{t}\left(\eta_{1},\eta_{2},\eta_{3},\omega_{1},\omega_{2},\omega_{3}\right)\\ \end{array}\right.$ (12)
 ${\mathrm{ARP}}_{t+1}=\frac{\hat{\mathrm{ARP}}_{t+1}}{\sqrt{\hat{\mathrm{ARP}}_{t+1}^{\prime}{\Sigma_{t}}C{\Sigma_{t}}\hat{\mathrm{ARP}}_{t+1}}}$ (13)
-3.4 Data and different simulated parameters
+4.4 Dataset and different simulated parameters
 The simulation starts on the 25th of may 1990 and stops on the 7th of december 2023.
@@
 5 Empirical results
-Here we first interpret the simulated empirical Sharpe ratio when applying the $\mathrm{ARP}$ portfolio with the diffent indicators. We find interesting results which challenge traditional recipes.
+Here we first interpret the simulated empirical Sharpe ratio when applying the $\mathrm{ARP}$ portfolio with the diffent indicators.
+We find interesting results which challenge traditional recipes.
+Indeed, and quite surprisingly, the empirical Sharpe ratio closely replicates the theoretical Sharpe ratio predicted by Grebenkov and Serror (2014), whose main result shows that a single $EMA$ indicator is already optimal.
+Secondly, we attempt to replicate the performance of a simple $EMA$ using a basket of more complex trend indicators.
+Thirdly, we incorporate the non-linear findings of Schmidhuber (2021), which could challenge our assumptions, and finally, we analyze the empirical correlations between different strategies.
 5.1 Grebenkov’s model empirical validation
+In Section 1 and 3, we introduced both the empirical and theoretical Sharpe ratios.
+As a reminder, the empirical Sharpe ratio is derived from backtesting, whereas the theoretical Sharpe ratio is obtained by modeling the underlying asset using theoretical diffusive processes.
+We tested Eq.
+3, which describes the theoretical Sharpe ratio as a function of the parameters of the trend indicator, as derived by Grebenkov and Serror (2014).
 Interestingly, the empirical simulation fits pretty well the theorical formula of the Sharpe ratio Eq.3 derived in Grebenkov and Serror (2014) with the following parameters $\lambda=1/180$ and $\beta_{0}=0.12$.
@@
 The theoritical model is fitted with parameters $\beta_{0}=0.12$ and $\lambda=1/180$ so that $\eta_{\text{opt}}=\frac{1}{112}$ Empirical measurements are displayed in the tab 4.
-5 Replication of a simple $\mathrm{EMA}$ by a mixture of bollinger bands $\mathrm{BB}$
+5.2 Replication of a simple $\mathrm{EMA}$ by a mixture of bollinger bands $\mathrm{BB}$
 The Bollinger Bands indicator, $\mathrm{BB}$, applies a double heavy-side function to an $\mathrm{SMA}$ with a width $\delta$.
@@
 Figure 6: weigths of a mixture of $\mathrm{BB}$ Vs size of the window to replicate the $\mathrm{EMA}\left(\frac{1}{112}\right)$ which is derived from Eq.16 corrected to pick different parameters of $\mathrm{BB}$ uniformely into lag in logarithm
+5.3 Non-linear findings which challenge the $EMA$
+Schmidhuber (2021) provides empirical evidence across different time scales suggesting that when trends are extreme, the optimal indicator becomes negatively dependent on the trend itself.
+As a consequence, strong trends appear to be relatively rare.
+Specifically, for a normalized trend variable $\phi$, they find that the optimal signal depends on a nonlinear transformation of the form $\phi-c\phi^{3}$, with $c=0.33$.
+They regress daily normalized returns on both the normalized trend and its cubic term as explanatory variables, across various underlyings and asset classes, including commodities.
+While they use a different trend indicator than $\phi$ derived in Eq.
+6, our version of the renormalized $\mathrm{EMA}$, their results challenge the assumption of a purely gaussian diffusive return process from Eq.
+1, which underpins the theoretical optimality of $\mathrm{EMA}$-based strategies.
+Interestingly, when applying the Agnostic Risk Parity (ARP) strategy using $\phi-c\phi^{3}$ instead of $\phi$ as the trend signal, we observe no improvement in the Sharpe ratio.
+This might be due to the lack of significance in the measurement of the coefficient $c$ (the sensitivity to the cubic term), which is highly sensitive to a few extreme trends, or to the loss of the rotational invariance property of the ARP (applying $\phi-c\phi^{3}$ to the raw returns or to their projections on the eigenvectors of the covariance matrix no longer yields equivalent results due to the nonlinearity of the transformation).
+A classical approach to introduce mean-reverting behavior from fundamentalist traders is to assume an attraction toward a fair or efficient price though such a price is unobservable.
+It can be approximated by a moving average, which effectively induces negative autocorrelation, especially over longer time scales.
+Zakamuli and Giner (2024) explores this idea, though obtaining statistically significant results remains difficult.
+By contrast, Schmidhuber (2021)’s approach involves making autoregressive coefficients trend-dependent, enabling a dynamic transition between trend-following governed by trend-followers and mean-reversion regimes governed by fundamentalist traders.
+Another way to model the influence of fundamental investors could be simply to assume their impact manifests in changes to the trend, i.e., on the noise $\xi_{i,k}$ in Eq.
+1.
+It could be enough to make sure that prices are not totally disconnected from fair and efficient ones which would differ in that case from moving average of prices.
+The findings of Schmidhuber (2021) should not be ignored.
+A deeper investigation is however warranted to better understand their implications for the construction of optimal trend-following portfolios, although this lies beyond the scope of the present simple work and should be the subject of a dedicated study.
 5.4 Correlations between indicators
-Fig.7 in the Appendix C shows that the indicators with different parameters give very correlated strategies which confims it makes not so appealing the taditional approach of using a basket of many indicators as signal.
+Fig.7 in the Appendix C shows that the indicators with different parameters give very correlated strategies which confims it makes not so appealing the traditional approach of using a basket of many indicators as signal.
 $\mathrm{ARP}\left(80\right)$ is correlated to $\mathrm{ARP}\left(150\right)$ with a coefficient of $0.96$.
@@
 So it makes this rafinement not justified again.
-5.2 Conclusion
+6 Conclusion
 Grebenkov’s model for describing trends is empirically validated, as its pretty complicated theoretical formula for determining the Sharpe ratio based on the $\mathrm{EMA}$ parameter fits impressively well with empirical data.
@@
 CHF/USD Fut.
-Table 2: List of the 41 instruments among currenies, equity indices and Bonds.
+Table 2: List of the 41 instruments among currencies, equity indices and Bonds.
 Appendix B Parameters

### v3

@@
 Abstract
-Our empirical results show an impressive fit with the pretty complex theoritical Sharpe formula of a Trend following strategy depending on the parameter of the signal, which was derived by Grebenkov and Serror (2014).
+Our empirical results show an impressive fit with the pretty complex theoretical Sharpe formula of a Trend following strategy depending on the parameter of the signal, which was derived by Grebenkov and Serror (2014).
 That empirical fit convinces us that a mean-reversion process with only one time scale is enough to model, in a pretty precise way, the reality of the trend-following mechanism at the average scale of CTAs and as a consequence, using only one simple $\mathrm{EMA}$, appears optimal to capture the trend.
 As a consequence, using a complex basket of different complex indicators as signal, do not seem to be so rational or optimal and exposes to the risk of cherry-picking.
-May 17, 2025
-I thank Prof Campbell Harvey, Robert Kosowsky, Valeriy Zakamulin, Denis Grebenkov and Raphael Douady for their feedbacks and valuable insights.
+August 7, 2025
+I thank Professors Campbell Harvey, Robert Kosowsky, Valeriy Zakamulin, Denis Grebenkov and Raphael Douady for their feedbacks and valuable insights.
 1 Introduction
-Practitioners of the trend-following strategies usually optimize, when replaying the historical scenarii through past returns, their strategies empirically by selecting parameters that maximize the simulated reward-to-risk ratio, typically measured by the Sharpe ratio.
+Practitioners of the trend-following strategies usually optimize their strategies empirically by selecting parameters that maximize the simulated reward-to-risk ratio, typically measured by the Sharpe ratio, when replaying the historical scenarii through past returns.
 However, this empirical approach can lead to overfitting, especially when optimization is performed on the same historical dataset used for evaluation.
@@
 It is also possible to determine the set of parameters that maximizes the theoretical Sharpe ratio—this is known as the theoretical optimal Sharpe ratio, as opposed to the common empirical optimum obtained through backtesting.
-A theoritical Sharpe ratio is always interesting if the model deriving the returns is realistic enough, as it allows to validate and optimize a strategy without the risk of overfiting.
+A theoretical Sharpe ratio is always interesting if the model deriving the returns is realistic enough, as it allows validating and optimizing a strategy without the risk of overfiting.
 The main objective of this paper is to empirically validate the theoretical Sharpe ratio model proposed by Grebenkov and Serror (2014) and therefore their assumption to model trends as gaussian mean reversion stochastic processes with only one time scale.
 2 Literature
-The usual recipe to determine a portfolio in the CTA’s industry is to use a blend of many different technical indicators as signal for each underlying instrument (most often individual indicator gives either a long or a short position).
-Then, a risk management process is applied to the signals to ensure that the portfolio is diversified enough and target a constant risk.
+The usual recipe to determine a portfolio in the CTA’s industry is to use a blend of many different technical indicators as a signal for each underlying instrument (most often individual indicator gives either a long or a short position111 Sepp and Lucic (2025) classified the methodologies for sizing positions based on signals into three categories: (1) European trend following, which assumes continuous position sizes proportional to signal strength — an approach used by many large European CTA managers; (2) Time Series Momentum (Moskowitz et al (2012); Hurst et al (2013); Baltas (2015)), a method more commonly used in academic studies, where positions depend on the sign of the momentum; and (3) American trend following, which originates from the early adopters of CTA strategies.).
+Then, a risk management process is applied to the signals to ensure that the portfolio is diversified enough and target s a constant risk.
 One simple solution is to impose an equal conditional risk in each asset class, as the SG Trend Indicator does and to size the portfolio to target a constant volatility.
@@
 In reality, one hidden assumption of Markowitz optimization is that the expected returns from the signals are known and certain, but this assumption is incorrect.
-Valeyre (2024) proved that the Agnostic Risk Parity ($\mathrm{ARP}$) approach was optimal when the correlation matrix between the trends or signals was a kind of random matrix where a dominant factor of this correlation matrix was very dominant and diffused randomnly.
+Valeyre (2024) proved that the Agnostic Risk Parity ($\mathrm{ARP}$) approach was optimal when the correlation matrix between the trends or signals was a kind of random matrix where a dominant factor of this correlation matrix was very dominant and diffused randomly.
 Valeyre (2024) suggested an interacting agents model to justify such an hypothesis.
@@
 The switch is determined when the short-term moving average crosses the long-term moving average.
-As an example, The Societe Generale Trend signal, which is a reference among CTAs, is detemined only by the crossover based on a simple average based 20 and 120 days parameters.
+As an example, The Societe Generale Trend signal, which is a reference among CTAs, is detemined only by the crossover based on a simple average based 20 and 120 day parameters.
 Momentum ($MOM$) is also a very popular indicator and is defined as the difference between two prices.
-Hurst et al (2017) analysed the performance simulated on the last century using a mixture of 1, 3 months and 1 year momentum.
-Lamperiere et al (2014) analysed the performance simulated on the last 2 centuries using a 50 days EMA signal.
-The Bollinger Band ($\mathrm{BB}$) is very popular but non continuous with only 3 outputs 0 and 1 or -1 when price is oustide a band.
-The Moving Average Convergence/Divergence ($\mathrm{MACD}$) uses a combination of three EMAs to capture mean reversion at both short- and long-term scales while identifying trends in the long term.
+Hurst et al (2017) analysed the performance simulated on the last century using a mixture of 1 month, 3 month and 1 year momentum.
+Lamperiere et al (2014) analysed the performance simulated on the last 2 centuries using a 50 day EMA signal.
+The Bollinger Band ($\mathrm{BB}$) is very popular, but non continuous with only 3 outputs 0 and 1 or -1 when price is oustide a band.
+The Moving Average Convergence/Divergence ($\mathrm{MACD}$) uses a combination of three EMAs to capture mean reversion at both short and long-term scales while identifying trends in the long term.
 We can describe some of these indicators through their sensitivity to past daily returns as introduced by Zakamuli and Giner (2020).
@@
 From an academic point of view, almost all these trend-following rules are ad hoc and lack theoretical justification for their optimality.
-Zakamuli and Giner (2020) related them to each other through the analysis of their sensitivity to past daily returns (so a 33 days EMA should be the optimal EMA to replicate the mixuture of 1, 3 months and 1 year momentum used by Hurst et al (2017) while a 43 days EMA should be the optimal one for the SG trend Indicator crossover).
+Zakamuli and Giner (2020) related them to each other through the analysis of their sensitivity to past daily returns (so a 33 day EMA should be the optimal EMA to replicate the mixuture of 1 month, 3 month and 1 year momentum used by Hurst et al (2017) while a 43 day EMA should be the optimal one for the SG trend Indicator crossover).
 Zakamuli and Giner (2020) also reminds us that it is well known that if returns follow an autoregressive process, the best predictor has the same functional form as the autoregressive process being predicted.
@@
 He also employs a two-state regime-switching model (bull and bear regimes), a widely accepted framework for stock returns (Timmermann (2000); Fruhwirth-Schnatter (2006); Giner and Zakamulin (2023)), using a semi-Markov model to detect negative autocorrelations over very long time scales.
-He argues that, within this framework, the MACD indicator with three different time scales could be adapted to construct the optimal portfolio but obtaining statistically significant results remains difficult.
+He argues that, within this framework, the MACD indicator with three different time scales could be adapted to construct the optimal portfolio, but obtaining statistically significant results remains difficult.
 So mean reversion at larger scales remains unconvincing.
-However the two-state regime does not seem natural for modeling trend-following mechanisms, which are better explained by herding behavior rather than by an external guiding force.
+However, the two-state regime does not seem natural for modeling trend-following mechanisms, which are better explained by herding behavior rather than by an external guiding force.
 Models in which agents partially imitate each other are more likely to produce continuous regime changes rather than abrupt shifts between two states.
@@
 This finding is in stark disagreement with Zakamuli and Giner (2024), simply because Grebenkov and Serror (2014) solved a different, yet superior, optimization problem.
-Indeed, assuming that positions depend linearly on the signal appears much more natural and is not as suboptimal as merely assuming a binary choice between a long or short position with the same level of risk withtout the possibility of implementing ”money management”.
+Indeed, assuming that positions depend linearly on the signal appears much more natural and is not as suboptimal as merely assuming a binary choice between a long or short position with the same level of risk withtout the possibility of implementing “money management”.
 Most important the version of Grebenkov and Serror (2014) is in line with Markowitz optimization and Agnostic risk parity by Benichou at al (2017) where positions are linearly depending on the signals.
@@
 Similarly, Nguyen at al (2014b) determined the optimal threshold for triggering either a long or short position.
-Our first objective was to validate empirically the model of Grebenkov and Serror (2014) through testing their beautiful formula describing the sensitivity of the empirical Sharpe ratio of the ’Agnostic Risk Portfolio’ ($\mathrm{ARP}$)—one of the components of the optimal trend-following portfolio derived by Valeyre (2024)—to the parameter of the $\mathrm{EMA}$.
-We also wanted to explore more complex signal than $\mathrm{EMA}$ to confirm that one time scale is enough to describe trends.
+Our first objective was to validate empirically the model of Grebenkov and Serror (2014) through testing their beautiful formula describing the sensitivity of the empirical Sharpe ratio of the “Agnostic Risk Portfolio” ($\mathrm{ARP}$)—one of the components of the optimal trend-following portfolio derived by Valeyre (2024)—to the parameter of the $\mathrm{EMA}$.
+We also wanted to explore more complex signals than $\mathrm{EMA}$ to confirm that one time scale is enough to describe trends.
+These empirical results challenge the common belief in our industry that complex systems are necessary.
+For example, Tzotchev (2018) derived an optimization scheme for assigning weights across different time scales using a HRP-Markowitz framework, based on theoretical correlations results derived from an AR(1) process for returns.
+It also challenges the very controversial  [^2] Buncic (2025); Elder (2025) empirical and theoretical findings of Kelly et al (2023), who theoretically argue that simple models severely understate return predictability.
 We decided to limit the other signals to a combination of three $\mathrm{EMA}$s to align with the $\mathrm{MACD}$ which could make sense if the autoregressive structure of returns could be more complex than the one which fits perfectly for a $\mathrm{EMA}$.
 However, there are two major differences:
-  • We aim to set the slope of the sensitivity at lag zero to zero as we focuse only on systematic returns through cross asset futures excluding specific risk.
+  • We aim to set the slope of the sensitivity at lag zero to zero as we focus only on systematic returns through cross asset futures excluding specific risk.
 We prefer to avoid a positive slope, as it would be curious for recent returns to have less impact on the trend than older ones and as short term Mean reversion behavior is known to be less important than for single stocks.
-  • We aim to achieve a larger fat tail for the larger scales instead of having a contrarian contribution which appears more adapted for capturing residual risk in single stocks but not for systematic part of returns.
+  • We aim to achieve a larger fat tail for the larger scales instead of having a contrarian contribution which appears more adapted at capturing residual risk in single stocks but not for systematic part of returns.
 Our second objective was to understand that a simple Exponential Moving Average ($\mathrm{EMA}$), which is supposed to be both optimal and simple, could also be replicated by a highly complex but more usual combination of indicators, such as Bollinger Bands, whose elementary indicator is even nonlinear and path-dependent.
@@
 Figure 1: Capture of Zakamuli and Giner (2024).
-”Spectrum” of different indicators based on Zakamuli and Giner (2024) i.e.
+“Spectrumm” of different indicators based on Zakamuli and Giner (2024) i.e.
 sensistivity to past daily returns.
@@
 A classical one-dimensional model used to capture trends in an underlying asset assumes that returns consist of two components: a random, independent and identically distributed (i.i.d.) component, and a small bias term.
-This bias can be modeled using an auto regressive process of order 1, AR(1), which represents mean reversion process around zero with only one time scale.
+This bias can be modeled using an auto-regressive process of order 1, AR(1), which represents mean reversion process around zero with only one time scale.
 The AR(1) process is characterized by one time scale parameter—indicating the average time needed to revert to zero—and a standard deviation that determines the level of noise.
@@
 The noises $\epsilon_{i,t}$ are independent in the $t$ axis but may be correlated with $i$ axis with a very close correlation matrix between returns.
-Through that model, the trend is $\beta\sum_{k=1}^{t-1}{\left(1-\lambda\right)^{t-1-k}\xi_{i,k}}$ and it follows an auto regressive model of order 1, AR(1).
+Through that model, the trend is $\beta\sum_{k=1}^{t-1}{\left(1-\lambda\right)^{t-1-k}\xi_{i,k}}$ and it follows an auto-regressive model of order 1, AR(1).
 $r_{i,t}=\epsilon_{i,t}+\beta\sum_{k=1}^{t-1}{\left(1-\lambda\right)^{t-1-k}\xi_{i,k}}$ (1)
+Note that an alternative model is also used in the literature, for example in Tzotchev (2018); Sepp and Lucic (2025), where the returns, rather than the trends, are modeled directly using an AR(1) process or even an autoregressive fractal process.
+However, this approach is less common than the classical model used by Grebenkov and Serror (2014).
 Through variogram measurements on the Dow Jones Index, they estimated that $\lambda=0.01$ and $\beta_{0}=0.1$ (Fig.2).
@@
 Another possible explanation is that the autocorrelation may not have been consistent over such a long period.
-Figure 2: Capture of Grebenkov and Serror (2014). Variogram applied to Dow Jones (1900-2012)to fit $\beta_{0}=0.08$ and $\lambda=0.011$ parameters
-Grebenkov and Serror (2014) solved the correct theoritical optimization problem—maximizing the Sharpe ratio at the portfolio level while ensuring that positions remain linearly dependent on signals when assuming that returns are well described by Eq.1.
+Figure 2: Capture of Grebenkov and Serror (2014).
+Variogram of standardized logarithmic daily returns (normalized by realized volatility) to Dow Jones (1900-2012) fit with $\beta_{0}=0.08$ and $\lambda=0.011$ parameters
+Grebenkov and Serror (2014) solved the correct theoretical optimization problem—maximizing the Sharpe ratio at the portfolio level while ensuring that positions remain linearly dependent on signals when assuming that returns are well described by Eq.1.
 They considered only the one-dimensional case, but later, Grebenkov and Serror (2015) and Valeyre (2024) extended the analysis to the multi-dimensional setting.
-Instead of focusing on the best predictor or on the optimal portfolio with suboptimal constraints—such as the possibility of being either long with the same risk or flat, which is unfortunately common in the literature—Grebenkov and Serror (2014) assumes that positions should better depend linearly on the signals, which is a natural approach which is in agreement with the Markowitz solution.
+Instead of focusing on the best predictor or on the optimal portfolio with suboptimal constraints—such as the possibility of being either long with the same risk or flat, which is unfortunately common in the literature—Grebenkov and Serror (2014) assumes that positions should better depend linearly on the signals [^3] They considered only the EMA as the signal, which was a good choice, as Zakamuli and Giner (2024) showed that the best predictor has the same functional form as the autoregressive process being predicted.
+That is, in the case of Eq.
+1, the optimal predictor should be an EMA, which they indeed used, which is a natural approach which is in agreement with the Markowitz solution.
 They then determine the optimal signal that maximizes the Sharpe ratio.
@@
 1) and $\beta$ (the weight of the trend component in the returns Eq.
-1) while also considering the inclusion of trading costs, denoted as $\theta$.
+1) or its normalized version $\beta_{0}$ while also considering the inclusion of trading costs, denoted as $\theta$.
 Equation Eq.3 is in fact a remarkably elegant formula.
-As expected, the theoretical Sharpe ratio depends on $\beta$, the weight of the trend in the returns, in an approximately proportional manner when $\beta_{0}^{2}$ is significantly larger than $\frac{1}{\lambda}$ and when $\eta$ remains close to $\lambda$.
+As expected, the theoretical Sharpe ratio depends on $\beta$ (or $\beta_{0}$ its normalized version), the weight of the trend in the returns, in an approximately proportional manner when $\beta_{0}^{2}$ is significantly larger than $\frac{1}{\lambda}$ and when $\eta$ remains close to $\lambda$.
 In that case, the Sharpe ratio from Eq.3 can be approximated by $\beta_{0}\sqrt{\frac{1}{2}}$ or $\beta_{0}\sqrt{\frac{255}{2}}$ in the annualized version of the Sharpe.
-Fig.3 illustrates the very complex theoritical formula Eq.3 through a simple graph which enable to undertsand its sensitivity against the parameter of the $\mathrm{EMA}$.
+Fig.3 illustrates the very complex theoretical formula Eq.3 through a simple graph that helps to understand its sensitivity to the $\mathrm{EMA}$ parameter..
 In reality, Eq.
@@
 Figure 3: Capture of Grebenkov and Serror (2014).
-Theoritical Sharpe (${\mathbf{S}}$ from Eq.
+Theoretical Sharpe ($\mathbf{S}$} from Eq.
 3) of the trend following strategy depending on $\eta$ the parameter of the $\mathrm{EMA}$.
@@
 This strategy constructs a rotationally invariant portfolio that applies trend-following while allocating equal unconditional risk to each eigenvector of the covariance matrix.
-While rotational invariance is a desirable theoretical property, the equal unconditional risk allocation among eigenvectors is a more desired property for asset manager.
+While rotational invariance is a desirable theoretical property, the equal unconditional risk allocation among eigenvectors is a more desired property for asset managers.
 Moreover, Valeyre (2024) demonstrated that the ARP portfolio is optimal—achieving the highest Sharpe ratio—under the assumption that the covariance of the $\xi_{i,k}$ terms in Eq.
@@
 The parameter $\sigma$ represents the standard deviation of returns used to normalize them, and $\eta$ is the smoothing parameter of the $EMA$ as well as of the return normalization.
-$\sigma^{2}$ can be interpreted as an exponential moving average of squared returns, and the indicator $\phi$ corresponds to an exponential moving average of returns scaled by their volatility..
-The incrementation is on daily basis but it could be generalized to minutes returns.
+$\sigma^{2}$ can be interpreted as an exponential moving average of squared returns, and the indicator $\phi$ corresponds to an exponential moving average of returns scaled by their volatility.
+The incrementation is on a daily basis but it could be generalized to minutes returns.
 $\sigma^{2}_{i,t+1}\left(\eta\right)=\left(1-\eta\right)\sigma^{2}_{i,t+1}\left(\eta\right)+\eta r^{2}_{i,t+1}$ (5)
-[…]
+$\phi_{i,t+1}\left(\eta\right)=\left(\ 1-\eta\right)\phi_{i,t}\left(\eta\right)+\sqrt{\eta}\frac{r_{i,t+1}}{\sigma_{i,t}\left(\eta\right)}$ (6)
 The indicator $\mathrm{EMA}\left(150\right)$ is defined by the $\phi_{i,t+1}\left(\eta\right)$ when $\eta=\frac{1}{150}$ is applied to every underlying $i$ and time $t$.
-We use $\sqrt{\eta}$ in Eq.6 so that the std of $\phi_{i,t+1}\left(\eta\right)$ is 1 theoritically if $r$ are well approximated by unautocorrelated returns so that $\phi$ is a normalized $EMA$ indicator.
+We use $\sqrt{\eta}$ in Eq.6 so that the std of $\phi_{i,t+1}\left(\eta\right)$ is 1 theoretically if $r$ are well approximated by unautocorrelated returns so that $\phi$ is a normalized $EMA$ indicator.
 4.2 Moving Average Convergence Divergence indicator ($\mathrm{MACD}$) as a 3 time scales $\mathrm{EMA}$ indicator
@@
 Finally, we applied a resizing process in Eq.13 to target a constant volatility for the final positions $\mathrm{ARP}$} which is a vector of N weights at time $t+1$.
-$\left\{\begin{array}[]{l}\hat{\mathrm{ARP}}_{t+1}=\left(1-\rho\right)\hat{\mathrm{ARP}}_{t}+\rho\Sigma^{-1}_{t}C^{-0.5}\phi_{t}\left(\eta\right)\\[10.0pt] or\\[10.0pt] \hat{\mathrm{ARP}}_{t+1}=\left(1-\rho\right)\hat{\mathrm{ARP}}_{t}+\rho\Sigma^{-1}_{t}C^{-0.5}\mathrm{MACD}_{t}\left(\eta_{1},\eta_{2},\eta_{3},\omega_{1},\omega_{2},\omega_{3}\right)\\ \end{array}\right.$ (12)
+[…]
 ${\mathrm{ARP}}_{t+1}=\frac{\hat{\mathrm{ARP}}_{t+1}}{\sqrt{\hat{\mathrm{ARP}}_{t+1}^{\prime}{\Sigma_{t}}C{\Sigma_{t}}\hat{\mathrm{ARP}}_{t+1}}}$ (13)
@@
 We tested the different indicators applying the $\mathrm{ARP}$ formula Eq.13. The different parameters are described in Tab 3 in the appendix B.
-Fig.4 displays the sensitivities of these indicators to past daily returns. $\mathrm{MACD}$ enables to put more weights on very older returns as we expected.
-Figure 4: Different sensitivities to past daily returns for indicators ($\mathrm{MACD}$, $\mathrm{EMA}$) all based on 20, 80 and 400 days time scales.
-we have $\mathrm{ARP}\left(0\times 20,80,0.2\right)$, $\mathrm{ARP}\left(0\times 20,80,0.4\times 400\right)$, $\mathrm{ARP}\left(20,80,0.2\times 400\right)$, $\mathrm{ARP}\left(20,80,0.4\times 400\right)$,$\mathrm{ARP}\left(80\right)$, $\mathrm{ARP}\left(400\right)$
+Fig.4 displays the sensitivities of these indicators to past daily returns. $\mathrm{MACD}$ enables us to put more weights on very older returns as we expected.
+Figure 4: Different sensitivities to past daily returns for indicators ($\mathrm{MACD}$, $\mathrm{EMA}$) all based on 20, 80 and 400 day time scales.
+we have $\mathrm{ARP}\left(0\times 20,80,0.2\right)$, $\mathrm{ARP}\left(0\times 20,80,0.4\times 400\right)$, $\mathrm{ARP}\left(20,80,0.2\times 400\right)$, $\mathrm{ARP}\left(20,80,0.4\times 400\right)$$\mathrm{ARP}\left(80\right)$, $\mathrm{ARP}\left(400\right)$
 5 Empirical results
-Here we first interpret the simulated empirical Sharpe ratio when applying the $\mathrm{ARP}$ portfolio with the diffent indicators.
+Here we first interpret the simulated empirical Sharpe ratio when applying the $\mathrm{ARP}$ portfolio with the different indicators.
 We find interesting results which challenge traditional recipes.
@@
 Interestingly, the empirical simulation fits pretty well the theorical formula of the Sharpe ratio Eq.3 derived in Grebenkov and Serror (2014) with the following parameters $\lambda=1/180$ and $\beta_{0}=0.12$.
-The Fig.5 is very impressive and the empricial fits should validate the Gebenkov’s model to describe trends.
-Moreover the fit appears to be more robust than the variogram (Fig.2) of the Dow Jones Index.
+The Fig.5 is very impressive and the empricial fits should validate Gebenkov’s model to describe trends.
+The risk of overfitting is very limited, as the formula was originally derived in 2014, with its parameters fitted solely on a long historical series of the Dow Jones Index.
+Using an extended dataset including commodities, forex, bonds, and other stock indices, we obtain parameter estimates that are consistent with those in Grebenkov and Serror (2014), even if they differ significantly in value.
+Moreover, the fit appears to be more robust than that obtained using the variogram method (Fig.
+2) applied to the Dow Jones Index, as our fitting is performed on a broader set of asset classes over a more recent period.
 That is the main result of the paper.
-As a first consequence the parameter of $112$ days ($\eta_{\text{opt}}=\frac{1}{112}$) for simple $\mathrm{EMA}$ is the optimal parameter to get the optimal Sharpe ratio when not accounting cost of trading which are very small at that trading frequency.
+As a first consequence the parameter of $112$ days ($\eta_{\text{opt}}=\frac{1}{112}$) for simple $\mathrm{EMA}$ is the optimal parameter to get the optimal Sharpe ratio when not accounting cost of trading which are very small at that trading frequency (Tab.
+4 displays holding periods slower than 80 days).
 Lamperiere et al (2014) found a different result, with a faster optimal parameter at 50 days for the EMA, but very similar Sharpe ratios across EMAs ranging from 1 to 5 months, showing an almost flat curve with no clear optimum.
@@
 For these two reasons, their Sharpe ratio for the post-2000 period was measured at 0.85, lower than our 1.2, and their optimal EMA signal corresponded to a faster timescale than ours.
-As a second consequence of the very good fit with the theoritical formula, modeling trend through mean-reversion process using only one relaxation time and not an multi-time scales one appears to be a good solution as the fit is more than correct.
-That is particulary unexpected as market is known to have a multi time scales property: For example the relaxation of volatility is known to have muti-time scales, investors are expected to have different horizons of time and different horizons of analysis.
-As a consequence we can wonder wether the usual recipe to take into account of a multitude of different indicators is justified to claim having the most robust and the most refined signal.
+As a second consequence of the very good fit with the theoretical formula, modeling trend through mean-reversion process using only one relaxation time and not an multi-time scales one appears to be a good solution as the fit is more than correct.
+That is particulary unexpected as market is known to have a multi-time scales property: For example the relaxation of volatility is known to have multi-time scales, investors are expected to have different horizons of time and different horizons of analysis.
+As a consequence we can wonder whether the usual recipe to take into account a multitude of different indicators is justified to claim having the most robust and the most refined signal.
+When properly renormalizing the signal, the position, and the portfolio, we do not observe trends reversing over longer horizons (e.g., 12 months) that would indicate mean reversion.
+Instead, the Grebenkov formula still holds even at horizons of up to 1,000 days.
+This stands in strong contradiction to the findings of Moskowitz et al (2012), who focused on momentum (MOM) indicators over the period 1965–2009.
 We can also note that $\beta_{0}=0.12$ is slightly higher than $0.08$, the parameter measured for the Dow Jones over the past 100 years by Grebenkov and Serror (2014).
@@
 We can see based on the Tab 4 in appendix C that empirical Sharpe is $1.24$ for $\mathrm{ARP}\left(120\right)$ with one time scale and $1.18$ for $\mathrm{MACD}\left(20,120,0.4\times 400\right)$ with 3 time scales.
-$\mathrm{MACD}$ could not be justified as additional time scales does not bring significant improvement.
+$\mathrm{MACD}$ could not be justified as additional time scales do not bring significant improvement.
 Also Sharpe ratio is not so sensitive to the parameters arround the optimal as expected: Sharpe is $1.25$ for $\mathrm{ARP}\left(100\right)$ and $1.21$ for $\mathrm{ARP}\left(150\right)$.
@@
 $\eta$ the parameter of the $\mathrm{EMA}$ (Eq.6) and the theoretical equation Eq.3.
-The theoritical model is fitted with parameters $\beta_{0}=0.12$ and $\lambda=1/180$ so that $\eta_{\text{opt}}=\frac{1}{112}$ Empirical measurements are displayed in the tab 4.
+The theoretical model is fitted with parameters $\beta_{0}=0.12$ and $\lambda=1/180$ so that $\eta_{\text{opt}}=\frac{1}{112}$ .
+Empirical measurements are displayed in the tab 4.
 5.2 Replication of a simple $\mathrm{EMA}$ by a mixture of Bollinger Bands $\mathrm{BB}$
@@
 $\mathrm{EMA}_{t}\left(\eta\right)=\frac{...+\left[\left(1-\eta\right)^{n-1}\eta\right]\left(n-1\right)\mathrm{SMA}_{t}\left(n-1\right)+...+\left[\left(1-\eta\right)\eta\right]\mathrm{SMA}_{t}\left(1\right)}{1+\left(1-\eta\right)+\left(1-\eta\right)^{2}+....+\left(1-\eta\right)^{n}+...}$ (15)
-[…]
+$\mathrm{SMA}_{t}\left(n\right)=\int_{0}^{\infty}\mathrm{BB}_{t}\left(n,\delta\right)\,d\delta$ (16)
 Figure 6: weigths of a mixture of $\mathrm{BB}$ Vs size of the window to replicate the $\mathrm{EMA}\left(\frac{1}{112}\right)$ which is derived from Eq.16 corrected to pick different parameters of $\mathrm{BB}$ uniformely into lag in logarithm
-5.3 Non-linear findings which challenge the $EMA$
+5.3 Non-linear findings which challenge the EMAEMA
 Schmidhuber (2021) provides empirical evidence across different time scales suggesting that when trends are extreme, the optimal indicator becomes negatively dependent on the trend itself.
@@
 It can be approximated by a moving average, which effectively induces negative autocorrelation, especially over longer time scales.
-Zakamuli and Giner (2024) explores this idea, though obtaining statistically significant results remains difficult.
+Zakamuli and Giner (2024) explore this idea, though obtaining statistically significant results remains difficult.
 By contrast, Schmidhuber (2021)’s approach involves making autoregressive coefficients trend-dependent, enabling a dynamic transition between trend-following governed by trend-followers and mean-reversion regimes governed by fundamentalist traders.
@@
 The findings of Schmidhuber (2021) should not be ignored.
-A deeper investigation is however warranted to better understand their implications for the construction of optimal trend-following portfolios, although this lies beyond the scope of the present simple work and should be the subject of a dedicated study.
+A deeper investigation is, however, warranted to better understand their implications for the construction of optimal trend-following portfolios, although this lies beyond the scope of the present simple work and should be the subject of a dedicated study.
 5.4 Correlations between indicators
-Fig.7 in the Appendix C shows that the indicators with different parameters give very correlated strategies which confims it makes not so appealing the traditional approach of using a basket of many indicators as signal.
+Fig.7 in Appendix C shows that the indicators with different parameters give very correlated strategies which confims that the traditional approach of using a basket of many indicators as signal is not so appealing.
 $\mathrm{ARP}\left(80\right)$ is correlated to $\mathrm{ARP}\left(150\right)$ with a coefficient of $0.96$.
-That is nevertheless pretty interesting to see so strong correlations which could be explained by a common factor somewhere, which could be the object of an additional research.
-Morevever the $\mathrm{ARP}\left(120\right)$ based on a simple $\mathrm{EMA}$ is very close with correlation from 1 to 0.99 to the new indicator $\mathrm{MACD}\left(20,120,0\times 400\right)$ or $\mathrm{MACD}\left(20,120,0.4\times 400\right)$ we introduced based on 3 different times scales to increase weights for very old past returns and decrease weights on very recent ones.
-So it makes this rafinement not justified again.
+That is nevertheless pretty interesting to see such strong correlations which could be explained by a common factor somewhere, which could be the object of an additional research.
+Morevever the $\mathrm{ARP}\left(120\right)$ based on a simple $\mathrm{EMA}$ is very close with correlation from 1 to 0.99 to the new indicator $\mathrm{MACD}\left(20,120,0\times 400\right)$ or $\mathrm{MACD}\left(20,120,0.4\times 400\right)$ we introduced based on 3 different time scales to increase weights for very old past returns and decrease weights on very recent ones.
+So it makes this refinement not justified again.
 6 Conclusion
@@
 The best fit is obtained using the theoretical model parameters $\lambda=1/180$ and $\beta_{0}=0.12$.
-As a consequence the parameter of $112$ days for simple $\mathrm{EMA}$ is the optimal parameter to get the optimal Sharpe ratio.
+As a consequence, the parameter of $112$ days for simple $\mathrm{EMA}$ is the optimal parameter to get the optimal Sharpe ratio.
 It is quite surprising that a single $\mathrm{EMA}$ is optimal for capturing trends, as one would expect different time scales for different types of investors.
@@
 The conclusion is that using a complex mixture of sophisticated indicators is unnecessary when the $\mathrm{EMA}$ alone provides a perfect fit—proving that simplicity can indeed be beautiful.
+7 Declaration of funding
+No funding was received
+8 Disclosure of interest
+There are no interests to declare,
 Appendix A Data

### v4

@@
 As a consequence, using a complex basket of different complex indicators as signal, do not seem to be so rational or optimal and exposes to the risk of cherry-picking.
-August 7, 2025
+October 24, 2025
 I thank Professors Campbell Harvey, Robert Kosowsky, Valeriy Zakamulin, Denis Grebenkov and Raphael Douady for their feedbacks and valuable insights.
@@
 2 Literature
-The usual recipe to determine a portfolio in the CTA’s industry is to use a blend of many different technical indicators as a signal for each underlying instrument (most often individual indicator gives either a long or a short position111 Sepp and Lucic (2025) classified the methodologies for sizing positions based on signals into three categories: (1) European trend following, which assumes continuous position sizes proportional to signal strength — an approach used by many large European CTA managers; (2) Time Series Momentum (Moskowitz et al (2012); Hurst et al (2013); Baltas (2015)), a method more commonly used in academic studies, where positions depend on the sign of the momentum; and (3) American trend following, which originates from the early adopters of CTA strategies.).
+The usual recipe to determine a portfolio in the CTA’s industry is to use a blend of many different technical indicators as a signal for each underlying instrument (most often individual indicator gives either a long or a short position [^1] Sepp and Lucic (2025) classified the methodologies for sizing positions based on signals into three categories: (1) European trend following, which assumes continuous position sizes proportional to signal strength — an approach used by many large European CTA managers; (2) Time Series Momentum (Moskowitz et al (2012); Hurst et al (2013); Baltas (2015)), a method more commonly used in academic studies, where positions depend on the sign of the momentum; and (3) American trend following, which originates from the early adopters of CTA strategies.).
 Then, a risk management process is applied to the signals to ensure that the portfolio is diversified enough and targets a constant risk.
@@
 The well-known Markowitz solution set positions as linearly depending on the signals.
-The linear dependence is derived simply througth the normalization of the signals by the inverse of the correlation matrix between returns, but the results are not appealing.
+The linear dependence is derived simply througth the normalization of the signals by the inverse of the correlation matrix between returns, but the empirical results are not appealing.
 The Agnostic Risk Parity ($\mathrm{ARP}$), introduced by Benichou at al (2017), normalizes the signals through the inverse of the square root of the correlation matrix.
@@
 Since this concept is primarily understood by physicists, their portfolio is not yet widely accepted in the finance community, with only a few citations, despite posing a serious challenge to Markowitz.
-In reality, one hidden assumption of Markowitz optimization is that the expected returns from the signals are known and certain, but this assumption is incorrect.
-Valeyre (2024) proved that the Agnostic Risk Parity ($\mathrm{ARP}$) approach was optimal when the correlation matrix between the trends or signals was a kind of random matrix where a dominant factor of this correlation matrix was very dominant and diffused randomly.
+For example, Benveniste at al (2024) explained, for the celebration of the 50th year of the Journal of Portfolio Management, the myth surrounding the alleged inadaptation of mean–variance optimization to the real world, and how efficient mean-variance optimization works without any conflict with empirical measurements, as long as the correlation matrix is properly modeled, yet neither referenced Benichou at al (2017) nor addressed the impact of uncertainty in mean within Markowitz formula.
+In reality, one hidden assumption of Markowitz optimization is that the expected means of returns from the signals are known and certain, but this assumption is incorrect.
+Valeyre (2024) proved that the Agnostic Risk Parity ($\mathrm{ARP}$) approach was optimal in the context of trend following, where expected means are not certain, are estimated with some errors and do not break rotational invariance as biases introduced by risk premia would, when the correlation matrix between the trends, signals or uncertainty of mean was a kind of random matrix where a dominant factor of this correlation matrix was very dominant and diffused randomly.
 Valeyre (2024) suggested an interacting agents model to justify such an hypothesis.
-Most of the popular technical indicators are based on moving averages of past prices.
+There are many different technical indicators used as proxies for expectated means in trend-following, but all of them are based on past returns, which are supposed to help predict future ones, while inevitably containing noise and uncertainty.
+Most of the popular indicators used are based on moving averages of past prices.
 The most popular is the Simple Moving Average ($\mathrm{SMA}$), while less commonly used types include the Linear Moving Average and the Exponential Moving Average on returns ($\mathrm{EMA}$).
@@
 From an academic point of view, almost all these trend-following rules are ad hoc and lack theoretical justification for their optimality.
-Zakamuli and Giner (2020) related them to each other through the analysis of their sensitivity to past daily returns (so a 33 day EMA should be the optimal EMA to replicate the mixuture of 1 month, 3 month and 1 year momentum used by Hurst et al (2017) while a 43 day EMA should be the optimal one for the SG trend Indicator crossover).
+Zakamuli and Giner (2020) related them to each other through the analysis of their sensitivity to past daily returns (so a 33 day EMA should be the optimal EMA to replicate the mixture of 1 month, 3 month and 1 year momentum used by Hurst et al (2017) while a 43 day EMA should be the optimal one for the SG trend Indicator crossover).
 Zakamuli and Giner (2020) also reminds us that it is well known that if returns follow an autoregressive process, the best predictor has the same functional form as the autoregressive process being predicted.
@@
 He argues that, within this framework, the MACD indicator with three different time scales could be adapted to construct the optimal portfolio, but obtaining statistically significant results remains difficult.
-So mean reversion at larger scales remains unconvincing.
+So mean reversion at larger scales remains in our opinion unconvincing despite the fact it was already documented by Moskowitz et al (2012).
 However, the two-state regime does not seem natural for modeling trend-following mechanisms, which are better explained by herding behavior rather than by an external guiding force.
@@
 A more credible representation would assume that trends follow an Ornstein-Uhlenbeck process or a similar mean-reverting process.
-Grebenkov and Serror (2014) proved that, in the more complex yet more insightful case where portfolio positions depend linearly on the signal, the $\mathrm{EMA}$ with a larger decay parameter than the decay parameter of the autoregressive coefficients of the return process yields the optimal portfolio.
+Grebenkov and Serror (2014) proved that, in the more complex yet more insightful case where portfolio positions depend linearly on the signal [^2] Grebenkov and Serror (2014) most likely introduced into the literature the ”linear trading rule” whose position depends linearly on the indicator, rather than a purely binary trading rule, which had been the standard approach most likely because the standard approach focuses on a one-dimensional problem and on a strategy that targets a constant volatility.
+We were unable to find any prior published articles, despite consulting ChatGPT, Google Scholar, and several professors in finance.
+However, as we are not specialist in this specific area of the literature, it is possible that we may have missed some references.
+Grebenkov and Serror (2014) introduced that linear methods in the literature because I was a colleague of both and I used that method to ensure that my statistical arbitrage trading program managing more than 500 dimensions or more than 500 single stocks at John Locke Investments maintained the property of “rotational invariance.” A linear trading rule was indeed necessary to have that rotational invariance property and design consistent strategies capable of performing trend following either on an individual stock, or a basket of stocks, or on the eigenvectors of the correlation matrix of individual stocks, rather than independently applying trend following to each stock in isolation.
+, the $\mathrm{EMA}$ with a larger decay parameter than the decay parameter of the autoregressive coefficients of the return process yields the optimal portfolio.
 This finding is in stark disagreement with Zakamuli and Giner (2024), simply because Grebenkov and Serror (2014) solved a different, yet superior, optimization problem.
 Indeed, assuming that positions depend linearly on the signal appears much more natural and is not as suboptimal as merely assuming a binary choice between a long or short position with the same level of risk withtout the possibility of implementing “money management”.
-Most important the version of Grebenkov and Serror (2014) is in line with Markowitz optimization and Agnostic risk parity by Benichou at al (2017) where positions are linearly depending on the signals.
-Moreover, Grebenkov and Serror (2014) is the only one to have derived an elegant and rigorous formula for the theoretical Sharpe ratio, explicitly linking it to the parameters of the return process and the relaxation time of the $\mathrm{EMA}$.
+Most important the version of Grebenkov and Serror (2014) is in line with the ”rotational invariance” property, the Markowitz optimization and Agnostic risk parity by Benichou at al (2017) where positions are linearly depending on the signals.
+The linear trading rule can be extended from the one-dimensional case to the multi-dimensional setting, allowing the incorporation of the correlation matrix to better adjust portfolio positions in order to target the maximum Sharpe ratio at the portfolio level.
+Moreover, Grebenkov and Serror (2014) is the only one to have derived an elegant333the formula proposed by Acar (2003), which depends on the correlation between the indicator and returns, leads-when expressed using the Grebenkov process-to an extremely cumbersome formula linking the parameters of the return process and the relaxation time of the EMA, which can hardly be described as objectively elegant.
+Furthermore, Zakamulin and Giner (2022) and Zakamuli and Giner (2024) are both based on the formula from Acar (2003) and rigorous formula for the theoretical Sharpe ratio, explicitly linking it to the parameters of the return process and the relaxation time of the $\mathrm{EMA}$.
 That formula is key in our main objective as we want to validate it using empirical measurements.
-Two years later, while Grebenkov and Serror (2014) remained largely unrecognized by the academic finance community, the literature continued to focus on the wrong optimization problem—one where positions do not depend linearly on the signals.
+Two years later, while Grebenkov and Serror (2014) remained largely unrecognized by the academic finance community 444the formula in Acar (2003); Zakamulin and Giner (2022); Zakamuli and Giner (2024) holds only with binary trading rules.
+Following Grebenkov and Serror (2014), Ferreira et al.
+(2018) published a similar formula for a linear trading rule, but in the context of a more general process with a simple moving average and not an EMA.
+Ferreira et al.
+(2018) did not mention Acar (2003) and derived his results through his own way.
+Later, after Grebenkov and Serror (2014), Firoozye and Koshiyama (2019) extended the formula of Acar (2003) to linear trading rules, yet these developments were ignored by Zakamulin and Giner (2022) and Zakamuli and Giner (2024), who continued to focuse exclusively on binary rules while ignoring Grebenkov and Serror (2014), the literature continued to focus on one-dimensional case and the wrong optimization problem—one where positions do not depend linearly on the signals.
 Dai at al (2016) addressed this issue by imposing a constraint that allowed only two possible positions (long or zero).
 Similarly, Nguyen at al (2014b) determined the optimal threshold for triggering either a long or short position.
-Our first objective was to validate empirically the model of Grebenkov and Serror (2014) through testing their beautiful formula describing the sensitivity of the empirical Sharpe ratio of the “Agnostic Risk Portfolio” ($\mathrm{ARP}$)—one of the components of the optimal trend-following portfolio derived by Valeyre (2024)—to the parameter of the $\mathrm{EMA}$.
-We also wanted to explore more complex signals than $\mathrm{EMA}$ to confirm that one time scale is enough to describe trends.
+Our first and main objective was to validate empirically the model of Grebenkov and Serror (2014) through testing their beautiful formula describing the sensitivity of the empirical Sharpe ratio of the “Agnostic Risk Portfolio” ($\mathrm{ARP}$)—one of the components of the optimal trend-following portfolio derived by Valeyre (2024)—to the parameter of the $\mathrm{EMA}$.
+This methodology ultimately extends the idea of Ferreira et al.
+(2018) by focusing on an AR(1) model to capture trends and by aggregating all markets through projection onto the $ARP$, rather than analyzing each market individually — which carries the risk of excessively noisy measurements and requires fitting complex and distinct ARIMA models.
+As a secondary result, we explored more complex signals than $\mathrm{EMA}$ to confirm that one time scale is enough to describe trends.
 These empirical results challenge the common belief in our industry that complex systems are necessary.
@@
 We decided to limit the other signals to a combination of three $\mathrm{EMA}$s to align with the $\mathrm{MACD}$ which could make sense if the autoregressive structure of returns could be more complex than the one which fits perfectly for a $\mathrm{EMA}$.
-However, there are two major differences:
-  • We aim to set the slope of the sensitivity at lag zero to zero as we focus only on systematic returns through cross asset futures excluding specific risk.
-We prefer to avoid a positive slope, as it would be curious for recent returns to have less impact on the trend than older ones and as short term Mean reversion behavior is known to be less important than for single stocks.
-  • We aim to achieve a larger fat tail for the larger scales instead of having a contrarian contribution which appears more adapted at capturing residual risk in single stocks but not for systematic part of returns.
-Our second objective was to understand that a simple Exponential Moving Average ($\mathrm{EMA}$), which is supposed to be both optimal and simple, could also be replicated by a highly complex but more usual combination of indicators, such as Bollinger Bands, whose elementary indicator is even nonlinear and path-dependent.
-This second objective raises the question: why use a complex combination of complex indicators that are sensitive to cherry-picking when a simple and elegant solution already exists?
+However, there are two main differences.
+First, we aim to set the slope of the sensitivity at lag zero to zero, as our focus is on systematic returns through cross-asset futures, excluding specific risk.
+A positive slope would be undesirable, since it would imply that recent returns have less influence on the trend than older ones, whereas short-term mean reversion behavior is known to be weaker for indices than for single stocks.
+Second, we seek to obtain a larger fat tail at longer scales rather than a contrarian contribution, which may be more appropriate for capturing residual risk in single stocks but not for the systematic component of returns.
+As another secondary result, we explored how a simple Exponential Moving Average ($\mathrm{EMA}$), which is supposed to be both optimal and simple, could also be replicated by a highly complex but more usual combination of indicators, such as Bollinger Bands, whose elementary indicator is even nonlinear and path-dependent.
+This second point raises the question: why use a complex combination of complex indicators that are sensitive to cherry-picking when a simple and elegant solution already exists?
 (a) MOM 100 days
@@
 Figure 1: Capture of Zakamuli and Giner (2024).
-“Spectrumm” of different indicators based on Zakamuli and Giner (2024) i.e.
+“Spectrum” of different indicators based on Zakamuli and Giner (2024) i.e.
 sensistivity to past daily returns.
@@
 In that case, the Sharpe ratio from Eq.3 can be approximated by $\beta_{0}\sqrt{\frac{1}{2}}$ or $\beta_{0}\sqrt{\frac{255}{2}}$ in the annualized version of the Sharpe.
-Fig.3 illustrates the very complex theoretical formula Eq.3 through a simple graph that helps to understand its sensitivity to the $\mathrm{EMA}$ parameter..
-In reality, Eq.
-3 was derived for the one-dimensional case only.
-Nevertheless, we can reasonably expect the equation to hold in the multi-dimensional setting.
-In that case, Eq.
+Fig.3 illustrates the very complex theoretical formula Eq.3 through a simple graph that helps to understand its sensitivity to the $\mathrm{EMA}$ parameter.
+[…]
+[…]
+Nevertheless, we can reasonably expect Eq. 3 to hold more precisely—thanks to rotational invariance—under the more realistic conditions that make the ARP, and not the Markowitz portfolio, optimal.
+In that case where Eq.
+3 holds, Eq.
 2 must be adjusted to redefine $\beta_{0}$.
-To be consistent with the observed increase of the Sharpe ratio with the size of the universe $N$, we can, under strong approximations, expect $\beta_{0}$ to be related to $\frac{\beta}{\sqrt{\lambda\left(2-\lambda\right)}}\sqrt{\frac{N}{1+\left(N-1\right)\rho^{2}}}$ where $\rho^{2}$ is the average squared correlation in the universe.
+To be consistent with the observed increase of the Sharpe ratio with the size of the universe $N$, we can, under strong approximations, expect $\beta_{0}$ to be related to $\frac{\beta}{\sqrt{\lambda\left(2-\lambda\right)}}\sqrt{\frac{N}{1+\left(N-1\right)\bm{\rho}^{2}}}$ where $\bm{\rho}^{2}$ is the average squared correlation in the universe.
 This expression extends Eq.
 2 using Eq.
-(40) from Grebenkov and Serror (2015).
-Additionally they derived in Eq.4, the optimal smoothing parameter ($\eta_{\text{opt}}$) of the Exponential Moving Average ($\mathrm{EMA}$) applied to returns as a trend indicator for a trend-following strategy which is supposed to yield to the optimal Sharpe ratio.
+(40)  [^10] Their Eq.
+(40) served only as an inspiration, as it was derived under the assumption of identical correlation coefficients for the two matrices (the correlations between $\xi_{i,k}$ and those between $\epsilon_{i,t}$)—conditions under which the Markowitz portfolio would be optimal, but not the $ARP$.
+This inconsistency is offset by another one: the Eq.
+(40) relates only the scaling factor between the “optimal Sharpe” and the “naïve Sharpe.”.
+from Grebenkov and Serror (2015).
+The rescaling factor formula $\sqrt{\frac{N}{1+\left(N-1\right)\bm{\rho}^{2}}}$ is, however, empirically tested through empirical Sharpe vs $N$ in Section 5.2.
+Additionally Grebenkov and Serror (2014) derived in Eq.4, the optimal smoothing parameter ($\eta_{\text{opt}}$) of the Exponential Moving Average ($\mathrm{EMA}$) applied to returns as a trend indicator for a trend-following strategy which is supposed to yield to the optimal Sharpe ratio.
 It is interesting to observe from Eq.4 that $\frac{1}{\eta_{\text{opt}}}$, the time scale of the signal $\mathrm{EMA}$ generating the optimal strategy is always shorter than $\frac{1}{\lambda}$, the time scale of the best predictor $\mathrm{EMA}$ (or the optimal indicator in the sub optimal optimization problem of Zakamuli and Giner (2024)) and that of the autocorrelated process of returns.
@@
 1 which is assumed for the underlying assets, we conducted empirical backtests of the Agnostic Risk Parity (ARP) strategy introduced by Benichou at al (2017).
-This strategy constructs a rotationally invariant portfolio that applies trend-following while allocating equal unconditional risk to each eigenvector of the covariance matrix.
+We decided to use the $ARP$ as the projection portfolio, as it is theoretically and empirically considered to be the optimal trend-following portfolio, capable of minimizing measurement noise in the Sharpe ratio which is empirically optimal (Benichou at al (2017)).
+However, we expect that the results should also hold even with lower Sharpe ratios and more noises for more conventional and naive trend-following portfolios, such as those that do not manage correlations or assume an identity correlation matrix except when computing volatility used to target the size of the portfolio.
+ARPARP constructs a rotationally invariant portfolio which makes the Sharpe ratio less dependent on the subjective selection of futures used to define the investment universe, as well as on discretionary choices regarding the relative number of FX, stock index, bond, and commodity futures.
+$ARP$ also by consequence applies trend-following to eigenvenvectors while allocating equal unconditional risk to each eigenvector of the covariance matrix.
 While rotational invariance is a desirable theoretical property, the equal unconditional risk allocation among eigenvectors is a more desired property for asset managers.
@@
 While the $EMA$ does not explicitly aim to avoid excessive trading, smoothing the portfolio has proven to be both simple and effective.
-Benichou at al (2017) recommended optimization under constraints to better manage trading costs).
+Benichou at al (2017) recommended optimization under constraints to limit trading and illiquidity and to better manage trading costs).
 Finally, we applied a resizing process in Eq.13 to target a constant volatility for the final positions $\mathrm{ARP}$} which is a vector of N weights at time $t+1$.
+The target-volatility step is both usual and useful to get standardized and homoscedastic returns for the portfolio so that Sharpe ratio is adapted.
+We applied exactly the same computation for the naïve approach, except that we assumed $C^{-0.5}$ in Eq.
+12 to be the identity matrix.
+This is equivalent to managing trend-following strategies independently, underlying by underlying, while constraining the overall portfolio to maintain constant volatility.
+This approach is referred to as ‘naïve’ by Benichou at al (2017).
+We did not test the true Markowitz portfolio, which would simply replace $C^{-0.5}$ by $C^{-1}$, as that solution performed significantly worse.
 […]
@@
 4.4 Dataset and different simulated parameters
-The simulation starts on the 25th of May 1990 and stops on the 7th of December 2023.
+The simulation starts on the 29th of May 1990 and stops on the 7th of December 2023.
 We used daily returns from 70 futures instruments in stock indices, bonds, fx and commodities futures.
@@
 3, which describes the theoretical Sharpe ratio as a function of the parameters of the trend indicator, as derived by Grebenkov and Serror (2014).
-Interestingly, the empirical simulation fits pretty well the theorical formula of the Sharpe ratio Eq.3 derived in Grebenkov and Serror (2014) with the following parameters $\lambda=1/180$ and $\beta_{0}=0.12$.
+Interestingly, the empirical simulation for ARP fits pretty well the theorical formula of the Sharpe ratio Eq.3 derived in Grebenkov and Serror (2014) with the following parameters $\lambda=1/180$ and $\beta_{0}=0.12$.
 The Fig.5 is very impressive and the empricial fits should validate Gebenkov’s model to describe trends.
-The risk of overfitting is very limited, as the formula was originally derived in 2014, with its parameters fitted solely on a long historical series of the Dow Jones Index.
-Using an extended dataset including commodities, forex, bonds, and other stock indices, we obtain parameter estimates that are consistent with those in Grebenkov and Serror (2014), even if they differ significantly in value.
+The risk of overfitting is very limited, and the fit was almost out-of sample as the formula was originally derived in 2014, with its parameters fitted solely on a long historical series of the Dow Jones Index.
+Using a 70 times larger dataset including commodities, forex, bonds, and other stock indices, we obtain parameter estimates that are consistent with those in Grebenkov and Serror (2014), even if they differ significantly in value.
 Moreover, the fit appears to be more robust than that obtained using the variogram method (Fig.
@@
 That is the main result of the paper.
+We can note that the measurements using $ARP$ fit the theoretical results better than those obtained with the naïve portfolio, where the correlation matrix is not considered to optimize diversification and reduce noise.
+The naïve measurements are not only shifted downward as we could have expected (the shift could have been more brutal if the universe was not well diversified to avoid potential criticism), but also appear slightly rotated (with $\lambda$ slightly increased) — a difference that may simply result from noisier estimates, or from the assumption made in Section 3 regarding the extension of Grebenkov’s formula, using the same $\lambda$, from the one-dimensional to the multi-dimensional case.
+It may be that Grebenkov’s formula holds more rigorously, from a mathematical standpoint, for the optimal $ARP$ portfolio than for the naïve one.
+However, this would be quite surprising, since Grebenkov’s formula is derived precisely under the mathematical conditions that make the naïve approach optimal (correlations between $\xi_{i,k}$ and those between $\epsilon_{i,t}$ are set to zero), as detailed in Section 3.
+The most likely explanation, which is not very intuitive since the error in the graph is highly autocorrelated with an estimated std of error of 0.18 in line with the 30 years of history, is that the $ARP$ measurements effectively apply a weighted average of $\lambda$ per market, assigning higher weights to markets which are different from other and which are not simply repetitions of others (as in the stock indices cases where all indices are very close).
+In contrast, the naïve approach uses an equal-weight scheme, which can overweight markets like stock indices that repeat themselves, thereby amplifying measurement errors.
+As a result, the naïve portfolio has just overweighted markets which appear faster.
+Schmidhuber (2021) similarly aggregates markets to obtain more robust results, observing that these coefficients are universal across asset classes and exhibit a universal scaling behavior, as the trend’s time horizon ranges from a few days to several years.
 As a first consequence the parameter of $112$ days ($\eta_{\text{opt}}=\frac{1}{112}$) for simple $\mathrm{EMA}$ is the optimal parameter to get the optimal Sharpe ratio when not accounting cost of trading which are very small at that trading frequency (Tab.
@@
 The difference with our results may be explained by the choice made in Lamperiere et al (2014) to determine positions based not on the linear magnitude of the EMA signals, but solely on their sign.
-Additionally, they did not use the ARP portfolio construction, which involves inverting the square root of the correlation matrix—a method that explains a significant part of the Sharpe ratio, as shown in Benichou at al (2017).
+Additionally, they use the ”naive” approach and did not use the ARP portfolio construction, which involves inverting the square root of the correlation matrix which slightly improves the fit, as observed in Fig.
+5.
 For these two reasons, their Sharpe ratio for the post-2000 period was measured at 0.85, lower than our 1.2, and their optimal EMA signal corresponded to a faster timescale than ours.
@@
 Also Sharpe ratio is not so sensitive to the parameters arround the optimal as expected: Sharpe is $1.25$ for $\mathrm{ARP}\left(100\right)$ and $1.21$ for $\mathrm{ARP}\left(150\right)$.
-Figure 5: Empirical Sharpe ratio based on the whole period 1990-2023 for trend following strategies using $\mathrm{ARP}$ (Eq.13 ) applied to $\mathrm{EMA}$ as signal v.s.
+Figure 5: Empirical Sharpe ratio based on the whole period 1990-2023 for trend following strategies using $\mathrm{ARP}$ (Eq.13 ) or the naive approach applied to $\mathrm{EMA}$ as signal v.s.
 $\eta$ the parameter of the $\mathrm{EMA}$ (Eq.6) and the theoretical equation Eq.3.
 The theoretical model is fitted with parameters $\beta_{0}=0.12$ and $\lambda=1/180$ so that $\eta_{\text{opt}}=\frac{1}{112}$ .
-Empirical measurements are displayed in the tab 4.
+Empirical measurements for the ARP are displayed in the tab 4.
+5.2 Scaling factor empirical validation
+We can see that the empirical measurements in Fig.
+5 appear less noisy than those reported by Ferreira et al.
+(2018) for the Dow Jones, where a complex ARMA process was required instead of a simple AR(1).
+It is clear that aggregating data by projecting onto our optimal trend-following portfolio, rather than analyzing each market individually, is highly effective in reducing noise that could otherwise create misleading evidence of mean reversion at longer time scales, as observed in Moskowitz et al (2012).
+We applied a simple empirical protocol to illustrate why analyzing the Sharpe ratio as a function of model parameters at the portfolio level is far more efficient than doing so market by market, as in Ferreira et al.
+(2018) (see their Figure 6), which was affected by considerable noise.
+Another advantage of this protocol is that it provides an empirical test of the scaling factor suggested in Section 3, where the Sharpe ratio depends on the size of the universe as $\sqrt{\frac{N}{1+\left(N-1\right)\rho^{2}}}$.
+We measured the empirical Sharpe ratios for the $ARP$ using an $EMA$ parameter value of $\frac{1}{120}$—close to the theoretical optimum—across several random universes containing 1, 3, 6, 9, 15, 20, and 27 underlying assets, each randomly selected from the original universe of 70 assets.
+We conducted 20 random trials for each case.
+Fig.
+6 shows another impressive fit, although the fitted value of $\bm{\rho}^{2}$ (0.024) appears slightly lower compared with the empirical average of squared correlations (0.056) based on the weekly returns.
+We can expect through that fit a Sharpe ratio of 1.28 for the whole universe of size 70, a Sharpe ratio of 1.40 for a 140 size and a Sharpe ratio of 1.60 for an infinite universe.
+The same figure also shows that the Sharpe ratio increases with the size of the universe, while its standard deviation, theoretically expected to remain stable around 0.2, declines only slightly—likely due to the larger correlations between randomly generated universes when they are larger.
+We can see that the mean Sharpe ratio is only about 0.2 when $N=1$, the one dimensional case, which is of the same order of magnitude as the noise level.
+This highlights how diversification and data aggregation—through projection onto the ARPARP—help reduce the noise-to-signal ratio of the measurements.
+Figure 6: Empirical Sharpe ratio based on the whole period 1990-2023 for trend following strategies using $\mathrm{ARP}$ (Eq.13 ) applied to $\mathrm{EMA}$ with $\eta=1/120$ as signal v.s.
+NN the size of the universe and the theoretical scaling factor inspired from Eq.
+(40) from Grebenkov and Serror (2015).
+The fitted $\bm{\rho}^{2}$ at 0.024 is just lower than the empirical average of squared correlations.
 5.3 Replication of a simple $\mathrm{EMA}$ by a mixture of Bollinger Bands $\mathrm{BB}$

### v5

@@
 As a consequence, using a complex basket of different complex indicators as signal, do not seem to be so rational or optimal and exposes to the risk of cherry-picking.
-October 24, 2025
+October 27, 2025
 I thank Professors Campbell Harvey, Robert Kosowsky, Valeriy Zakamulin, Denis Grebenkov and Raphael Douady for their feedbacks and valuable insights.
@@
 from Grebenkov and Serror (2015).
-The rescaling factor formula $\sqrt{\frac{N}{1+\left(N-1\right)\bm{\rho}^{2}}}$ is, however, empirically tested through empirical Sharpe vs $N$ in Section 5.2.
+The rescaling factor formula, $\sqrt{\frac{N}{1+(N-1)\bm{\rho}^{2}}}$, is empirically tested through the relationship between the Sharpe ratio and $N$ in Section 5.2.
+This rescaling factor formula also corresponds exactly to the theoretical scaling of the Sharpe ratio for a portfolio of $N$ assets with identical expected returns and volatilities, and an average pairwise correlation of $\rho^{2}$.
+This makes sense, as the correlation between two trend-following strategies built on underlyings correlated at $\rho$ is expected to be $\rho^{2}$.
 Additionally Grebenkov and Serror (2014) derived in Eq.4, the optimal smoothing parameter ($\eta_{\text{opt}}$) of the Exponential Moving Average ($\mathrm{EMA}$) applied to returns as a trend indicator for a trend-following strategy which is supposed to yield to the optimal Sharpe ratio.
@@
 The target-volatility step is both usual and useful to get standardized and homoscedastic returns for the portfolio so that Sharpe ratio is adapted.
-We applied exactly the same computation for the naïve approach, except that we assumed $C^{-0.5}$ in Eq.
+We applied exactly the same computation for the naive approach, except that we assumed $C^{-0.5}$ in Eq.
 12 to be the identity matrix.
@@
 We did not test the true Markowitz portfolio, which would simply replace $C^{-0.5}$ by $C^{-1}$, as that solution performed significantly worse.
+We also test the naive approach derived from the binary trading rule where we replace $\phi_{t}$ by its sign.
 […]
@@
 However, this would be quite surprising, since Grebenkov’s formula is derived precisely under the mathematical conditions that make the naïve approach optimal (correlations between $\xi_{i,k}$ and those between $\epsilon_{i,t}$ are set to zero), as detailed in Section 3.
-The most likely explanation, which is not very intuitive since the error in the graph is highly autocorrelated with an estimated std of error of 0.18 in line with the 30 years of history, is that the $ARP$ measurements effectively apply a weighted average of $\lambda$ per market, assigning higher weights to markets which are different from other and which are not simply repetitions of others (as in the stock indices cases where all indices are very close).
-In contrast, the naïve approach uses an equal-weight scheme, which can overweight markets like stock indices that repeat themselves, thereby amplifying measurement errors.
+The most likely explanation, which is not very intuitive since the error in the graph is highly autocorrelated with an estimated std of error of 0.18 in line with the 30 years of history, is that the $ARP$ measurements effectively apply a similar process to the weighted average of $\lambda$ per market, assigning higher weights to markets which are different from other and which are not simply repetitions of others (as in the stock indices cases where all indices are very close).
+In contrast, the naïve approach uses a similar process to an equal-weight scheme, which can overweight markets like stock indices that repeat themselves, thereby amplifying measurement errors.
 As a result, the naïve portfolio has just overweighted markets which appear faster.
 Schmidhuber (2021) similarly aggregates markets to obtain more robust results, observing that these coefficients are universal across asset classes and exhibit a universal scaling behavior, as the trend’s time horizon ranges from a few days to several years.
+We can also note that the $ARP$ approach yields higher Sharpe ratios than the naïve approach, even when binary trading rules are used, the overperformance would have been a lot larger if the universe was not as diversified.
 As a first consequence the parameter of $112$ days ($\eta_{\text{opt}}=\frac{1}{112}$) for simple $\mathrm{EMA}$ is the optimal parameter to get the optimal Sharpe ratio when not accounting cost of trading which are very small at that trading frequency (Tab.
@@
 Also Sharpe ratio is not so sensitive to the parameters arround the optimal as expected: Sharpe is $1.25$ for $\mathrm{ARP}\left(100\right)$ and $1.21$ for $\mathrm{ARP}\left(150\right)$.
-Figure 5: Empirical Sharpe ratio based on the whole period 1990-2023 for trend following strategies using $\mathrm{ARP}$ (Eq.13 ) or the naive approach applied to $\mathrm{EMA}$ as signal v.s.
+Figure 5: Empirical Sharpe ratio based on the whole period 1990-2023 for trend following strategies using $\mathrm{ARP}$ (Eq.13 ) or the naive approach applied to $\mathrm{EMA}$ as signal (both with the linear trading rule and the binary trading rules) v.s.
 $\eta$ the parameter of the $\mathrm{EMA}$ (Eq.6) and the theoretical equation Eq.3.
@@
 We can expect through that fit a Sharpe ratio of 1.28 for the whole universe of size 70, a Sharpe ratio of 1.40 for a 140 size and a Sharpe ratio of 1.60 for an infinite universe.
+We can also observe that the fit obtained with the simpler square-root formula is less accurate and would predict larger Sharpe ratios for larger universes.
 The same figure also shows that the Sharpe ratio increases with the size of the universe, while its standard deviation, theoretically expected to remain stable around 0.2, declines only slightly—likely due to the larger correlations between randomly generated universes when they are larger.
@@
 This highlights how diversification and data aggregation—through projection onto the ARPARP—help reduce the noise-to-signal ratio of the measurements.
+We can also see that the exact standard deviation at $N=1$ of 0.2087 is not significantly larger than the standard deviation of 0.175, just explained by the noise of Sharpe ratio using 32.58 years of data, that we would also obtain with an infinite number of random trials instead of 20, assuming that the autocorrelation coefficients were universal across asset classes.
+This last observation justifies our choice to aggregate data across asset classes.
 Figure 6: Empirical Sharpe ratio based on the whole period 1990-2023 for trend following strategies using $\mathrm{ARP}$ (Eq.13 ) applied to $\mathrm{EMA}$ with $\eta=1/120$ as signal v.s.

### v6

@@
 As a consequence, using a complex basket of different complex indicators as signal, do not seem to be so rational or optimal and exposes to the risk of cherry-picking.
-October 27, 2025
+November 3, 2025
 I thank Professors Campbell Harvey, Robert Kosowsky, Valeriy Zakamulin, Denis Grebenkov and Raphael Douady for their feedbacks and valuable insights.
@@
 However, this empirical approach can lead to overfitting, especially when optimization is performed on the same historical dataset used for evaluation.
+Many financial researchers have proposed various methods to address this issue, which can be broadly classified into three categories, as noted by Koshiyama and Firoozye (2019): data snooping, overestimated performance, and cross-validation evaluation.
+Koshiyama and Firoozye (2019) introduce a technique to mitigate financial overfitting called the Covariance-Penalty Correction, which adjusts a risk metric downward based on the number of parameters and the volume of data underpinning a trading strategy.
 An alternative is to model the returns of the underlying assets using stochastic processes.
@@
 (f) MACD 30-100-400 days
-Figure 1: Capture of Zakamuli and Giner (2024).
-“Spectrum” of different indicators based on Zakamuli and Giner (2024) i.e.
+Figure 1: “Spectrum” of different indicators inspired by Zakamuli and Giner (2024) i.e.
 sensistivity to past daily returns.
@@
 Another possible explanation is that the autocorrelation may not have been consistent over such a long period.
-Figure 2: Capture of Grebenkov and Serror (2014).
-Variogram of standardized logarithmic daily returns (normalized by realized volatility) to Dow Jones (1900-2012) fit with $\beta_{0}=0.08$ and $\lambda=0.011$ parameters
+Figure 2: Variogram of standardized logarithmic daily returns (normalized by realized volatility) to Dow Jones (1900-2012) fit with $\beta_{0}=0.08$ and $\lambda=0.011$ parameters, replicated from Grebenkov and Serror (2014)
 Grebenkov and Serror (2014) solved the correct theoretical optimization problem—maximizing the Sharpe ratio at the portfolio level while ensuring that positions remain linearly dependent on signals when assuming that returns are well described by Eq.1.
@@
 Fig.3 illustrates the very complex theoretical formula Eq.3 through a simple graph that helps to understand its sensitivity to the $\mathrm{EMA}$ parameter.
-[…]
-[…]
-Nevertheless, we can reasonably expect Eq. 3 to hold more precisely—thanks to rotational invariance—under the more realistic conditions that make the ARP, and not the Markowitz portfolio, optimal.
+In reality, Eq.
+3 was derived for the one-dimensional case only but it could be easily extended to the multi-dimensional case: Obviously, Grebenkov and Serror (2015) proved that Eq.
+3 holds when the correlations between $\xi_{i,k}$ and those between $\epsilon_{i,t}$ are set to zero, as shown in their Eq.
+(24) [^7] Eq.
+(24) $\sqrt{\frac{Nq^{2}(1-p^{2})}{Q^{2}+2Q+R}}$ where $p=1-\eta$, $q=1-\lambda$, $Q=\frac{(1-pq)\sigma^{4}}{\beta_{0}^{4}}$, and $R=1-q^{2}-2p^{2}q^{2}$..
+Eq.
+3 also holds when the correlations between $\xi_{i,k}$ are the same as those between $\epsilon_{i,t}$, as shown in their Eq.
+(39) [^8] Eq.
+(39) $\sqrt{\frac{Nq^{2}(1-p^{2})}{Q^{2}+2Q+R}}$ where $p=1-\eta$, $q=1-\lambda$, $Q=\frac{(1-pq)\sigma^{4}}{\beta_{0}^{4}}$, and $R=1-q^{2}-2p^{2}q^{2}$., with both cases involving a scaling factor that depends only on the number of underlying assets.
+These latter conditions also make the Markowitz portfolio optimal, as proved by Valeyre (2024), but they are not realistic: the resulting “optimal Sharpe” would increase linearly with the square root of the number of underlyings, with the same slope regardless of correlation—an outcome that would be too ideal and profitable to be credible.
+This in practice, with large $N$ yields an expected Sharpe ratio based on in-sample data that is much higher—by several orders of magnitude—than the out-of-sample Sharpe ratio.
+This is the signature of overfitting, which occurs simply because the model assumes that the estimated means are measured without error especially for small eigenvectors, thereby allowing unrealistically profitable hedges to reduce risk under the illusion of certainty about expected mean of returns.
+Eq.
+3 differs slightly when assuming that the correlations between $\xi_{i,k}$ are zero, as shown in their Eq.
+(43) [^9] Eq.
+(43) $\sqrt{\frac{Nq^{2}(1-p^{2})}{Q^{2}(1-\bm{\rho_{\epsilon}}^{2})+2Q(1-\bm{\rho_{\epsilon}})+R}}$ where $p=1-\eta$, $q=1-\lambda$, and $Q=\frac{(1-pq)\sigma^{4}}{\beta_{0}^{4}}$., but this correlation-dependent formulation remains problematic: it fails to converge as $N\to\infty$, which contradicts empirical observations, suggesting that the underlying assumption (that correlations between $\xi_{i,k}$ are zero) is unrealistic.
+Nevertheless, we can reasonably expect Eq.
+3 to hold more precisely and perhaps even exactly—thanks to rotational invariance  [^10] We know from Grebenkov and Serror (2015) and the discussion above that Eq.
+3 holds when considering coordinate transformation to the eigenvectors of the correlation matrix of $\epsilon_{i,t}$, in which the ”random” correlations between $\xi_{i,k}$ and those between $\epsilon_{i,t}$ would be set to zero under the $ARP$ conditions.—under the more realistic conditions that make the ARP, and not the Markowitz portfolio, optimal.
 In that case where Eq.
@@
 This inconsistency is offset by another one: the Eq.
-(40) relates only the scaling factor between the “optimal Sharpe” and the “naïve Sharpe.”.
+(40) relates only the scaling factor between the optimal Sharpe and the naïve ’1/N’ Sharpe.
 from Grebenkov and Serror (2015).
@@
 This makes sense, as the correlation between two trend-following strategies built on underlyings correlated at $\rho$ is expected to be $\rho^{2}$.
+Thus, the scaling factor for ARP ultimately coincides with that obtained for the naïve ’1/N’ trend-following portfolio, which does not use the correlation matrix to determine positions.
 Additionally Grebenkov and Serror (2014) derived in Eq.4, the optimal smoothing parameter ($\eta_{\text{opt}}$) of the Exponential Moving Average ($\mathrm{EMA}$) applied to returns as a trend indicator for a trend-following strategy which is supposed to yield to the optimal Sharpe ratio.
@@
 $\eta_{\text{opt}}=\lambda\sqrt{1+2\frac{\beta^{2}_{0}}{\lambda}}$ (4)
-Figure 3: Capture of Grebenkov and Serror (2014).
-Theoretical Sharpe ($\mathbf{S}$} from Eq.
+Figure 3: Theoretical Sharpe ($\mathbf{S}$} from Eq.
 3) of the trend following strategy depending on $\eta$ the parameter of the $\mathrm{EMA}$.
-$\beta_{0}=0.1$ and $\lambda=0.01$ for different cost of trading $\theta$
+$\beta_{0}=0.1$ and $\lambda=0.01$ for different cost of trading $\theta$, replicated from Grebenkov and Serror (2014)
 4 Description of the empirical analysis
@@
 We decided to use the $ARP$ as the projection portfolio, as it is theoretically and empirically considered to be the optimal trend-following portfolio, capable of minimizing measurement noise in the Sharpe ratio which is empirically optimal (Benichou at al (2017)).
-However, we expect that the results should also hold even with lower Sharpe ratios and more noises for more conventional and naive trend-following portfolios, such as those that do not manage correlations or assume an identity correlation matrix except when computing volatility used to target the size of the portfolio.
+However, we expect that the results should also hold even with lower Sharpe ratios and more noises for more conventional and naive ’1/N’ trend-following portfolios, such as those that do not manage correlations or assume an identity correlation matrix except when computing volatility used to target the size of the portfolio.
 ARPARP constructs a rotationally invariant portfolio which makes the Sharpe ratio less dependent on the subjective selection of futures used to define the investment universe, as well as on discretionary choices regarding the relative number of FX, stock index, bond, and commodity futures.
@@
 $0=\omega_{1}\sqrt{\eta_{1}}+\omega_{2}\sqrt{\eta_{2}}+\omega_{3}\sqrt{\eta_{3}}$ (8)
-4.3 Agnostic Risk Parity ($\mathrm{ARP}$) when smoothing the portfolio and targeting a constant volatility
+4.3 Agnostic Risk Parity ($\mathrm{ARP}$) and naive ’1/N’ portfolio
 We first estimated the correlation matrix $C$, of dimension $N\times N$, using a 750-day exponential moving average applied to weekly returns when implemnting the RIE filter introduced by Bun at al (2016).
@@
 The target-volatility step is both usual and useful to get standardized and homoscedastic returns for the portfolio so that Sharpe ratio is adapted.
-We applied exactly the same computation for the naive approach, except that we assumed $C^{-0.5}$ in Eq.
+We applied exactly the same computation for the naive ’1/N’ approach, except that we assumed $C^{-0.5}$ in Eq.
 12 to be the identity matrix.
 This is equivalent to managing trend-following strategies independently, underlying by underlying, while constraining the overall portfolio to maintain constant volatility.
-This approach is referred to as ‘naïve’ by Benichou at al (2017).
+This naive approach is referred to as ‘1/N’ by Benichou at al (2017).
 We did not test the true Markowitz portfolio, which would simply replace $C^{-0.5}$ by $C^{-1}$, as that solution performed significantly worse.
-We also test the naive approach derived from the binary trading rule where we replace $\phi_{t}$ by its sign.
+We also test the naive ’1/N’ approach derived from the binary trading rule where we replace $\phi_{t}$ by its sign.
 […]
@@
 3, which describes the theoretical Sharpe ratio as a function of the parameters of the trend indicator, as derived by Grebenkov and Serror (2014).
-Interestingly, the empirical simulation for ARP fits pretty well the theorical formula of the Sharpe ratio Eq.3 derived in Grebenkov and Serror (2014) with the following parameters $\lambda=1/180$ and $\beta_{0}=0.12$.
-The Fig.5 is very impressive and the empricial fits should validate Gebenkov’s model to describe trends.
-The risk of overfitting is very limited, and the fit was almost out-of sample as the formula was originally derived in 2014, with its parameters fitted solely on a long historical series of the Dow Jones Index.
-Using a 70 times larger dataset including commodities, forex, bonds, and other stock indices, we obtain parameter estimates that are consistent with those in Grebenkov and Serror (2014), even if they differ significantly in value.
-Moreover, the fit appears to be more robust than that obtained using the variogram method (Fig.
-2) applied to the Dow Jones Index, as our fitting is performed on a broader set of asset classes over a more recent period.
-That is the main result of the paper.
-We can note that the measurements using $ARP$ fit the theoretical results better than those obtained with the naïve portfolio, where the correlation matrix is not considered to optimize diversification and reduce noise.
-The naïve measurements are not only shifted downward as we could have expected (the shift could have been more brutal if the universe was not well diversified to avoid potential criticism), but also appear slightly rotated (with $\lambda$ slightly increased) — a difference that may simply result from noisier estimates, or from the assumption made in Section 3 regarding the extension of Grebenkov’s formula, using the same $\lambda$, from the one-dimensional to the multi-dimensional case.
-It may be that Grebenkov’s formula holds more rigorously, from a mathematical standpoint, for the optimal $ARP$ portfolio than for the naïve one.
-However, this would be quite surprising, since Grebenkov’s formula is derived precisely under the mathematical conditions that make the naïve approach optimal (correlations between $\xi_{i,k}$ and those between $\epsilon_{i,t}$ are set to zero), as detailed in Section 3.
+[…]
+We can note that the measurements using $ARP$ fit the theoretical results better than those obtained with the naïve ’1/N’ portfolio (using the scipy.optimize library, we obtained for the naive ’1/N’ case $\lambda=1/110$ but $R^{2}=0.75$), where the correlation matrix is not considered to optimize diversification and reduce noise.
+The naïve ’1/N’ measurements are not only shifted downward as we could have expected (the shift could have been more brutal if the universe was not well diversified to avoid potential criticism), but also appear slightly rotated (with $\lambda$ slightly increased) — a difference that may simply result from noisier estimates, or from the assumption made in Section 3 regarding the extension of Grebenkov’s formula, using the same $\lambda$, from the one-dimensional to the multi-dimensional case.
+It may be that Grebenkov’s formula holds more rigorously, from a mathematical standpoint, for the optimal $ARP$ portfolio than for the naïve ’1/N’ one.
+However, this would be quite surprising, since Grebenkov’s formula is derived precisely under the mathematical conditions that make the naïve ’1/N’ approach optimal (correlations between $\xi_{i,k}$ and those between $\epsilon_{i,t}$ are set to zero), as detailed in Section 3.
 The most likely explanation, which is not very intuitive since the error in the graph is highly autocorrelated with an estimated std of error of 0.18 in line with the 30 years of history, is that the $ARP$ measurements effectively apply a similar process to the weighted average of $\lambda$ per market, assigning higher weights to markets which are different from other and which are not simply repetitions of others (as in the stock indices cases where all indices are very close).
-In contrast, the naïve approach uses a similar process to an equal-weight scheme, which can overweight markets like stock indices that repeat themselves, thereby amplifying measurement errors.
-As a result, the naïve portfolio has just overweighted markets which appear faster.
+In contrast, the naïve ’1/N’ approach uses a similar process to an equal-weight scheme, which can overweight markets like stock indices that repeat themselves, thereby amplifying measurement errors.
+As a result, the naïve ’1/N’ portfolio has just overweighted markets which appear faster.
 Schmidhuber (2021) similarly aggregates markets to obtain more robust results, observing that these coefficients are universal across asset classes and exhibit a universal scaling behavior, as the trend’s time horizon ranges from a few days to several years.
-We can also note that the $ARP$ approach yields higher Sharpe ratios than the naïve approach, even when binary trading rules are used, the overperformance would have been a lot larger if the universe was not as diversified.
+We can also note that the $ARP$ approach yields higher Sharpe ratios than the naïve ’1/N’ approach, even when binary trading rules are used, the overperformance would have been a lot larger if the universe was not as diversified.
 As a first consequence the parameter of $112$ days ($\eta_{\text{opt}}=\frac{1}{112}$) for simple $\mathrm{EMA}$ is the optimal parameter to get the optimal Sharpe ratio when not accounting cost of trading which are very small at that trading frequency (Tab.
@@
 The difference with our results may be explained by the choice made in Lamperiere et al (2014) to determine positions based not on the linear magnitude of the EMA signals, but solely on their sign.
-Additionally, they use the ”naive” approach and did not use the ARP portfolio construction, which involves inverting the square root of the correlation matrix which slightly improves the fit, as observed in Fig.
+Additionally, they use the naive ’1/N’ approach and did not use the ARP portfolio construction, which involves inverting the square root of the correlation matrix which slightly improves the fit, as observed in Fig.
 5.
@@
 This further confirms the importance of measuring the implied autocorrelation parameter based on a strategy invested in a large universe, as it leads to more accurate and agregated estimates.
-We can also note that $\lambda$ is estimated to $\frac{1}{180}$ instead of $0.011$ in the case of the Dow Jones.
+We can also note that $\lambda$ is estimated to $\frac{1}{180\pm 17}$ instead of $0.011$ in the case of the Dow Jones.
 Our analysis is that market behavior should have changed in the last 100 years and we believe that our fit appears more robust than a simple variogram.
@@
 Also Sharpe ratio is not so sensitive to the parameters arround the optimal as expected: Sharpe is $1.25$ for $\mathrm{ARP}\left(100\right)$ and $1.21$ for $\mathrm{ARP}\left(150\right)$.
-Figure 5: Empirical Sharpe ratio based on the whole period 1990-2023 for trend following strategies using $\mathrm{ARP}$ (Eq.13 ) or the naive approach applied to $\mathrm{EMA}$ as signal (both with the linear trading rule and the binary trading rules) v.s.
+Figure 5: Empirical Sharpe ratio based on the whole period 1990-2023 for trend following strategies using $\mathrm{ARP}$ (Eq.13 ) or the naive ’1/N’ approach applied to $\mathrm{EMA}$ as signal (both with the linear trading rule and the binary trading rules) v.s.
 $\eta$ the parameter of the $\mathrm{EMA}$ (Eq.6) and the theoretical equation Eq.3.
@@
 Fig.
-6 shows another impressive fit, although the fitted value of $\bm{\rho}^{2}$ (0.024) appears slightly lower compared with the empirical average of squared correlations (0.056) based on the weekly returns.
+6 shows another impressive fit, although the fitted value of $\bm{\rho}^{2}$ ($0.024\pm 0.012$) appears slightly lower compared with the empirical average of squared correlations ($0.056$) based on the weekly returns.
 We can expect through that fit a Sharpe ratio of 1.28 for the whole universe of size 70, a Sharpe ratio of 1.40 for a 140 size and a Sharpe ratio of 1.60 for an infinite universe.
 We can also observe that the fit obtained with the simpler square-root formula is less accurate and would predict larger Sharpe ratios for larger universes.
-The same figure also shows that the Sharpe ratio increases with the size of the universe, while its standard deviation, theoretically expected to remain stable around 0.2, declines only slightly—likely due to the larger correlations between randomly generated universes when they are larger.
+The same Fig.
+6 also shows that the Sharpe ratio increases with the size of the universe, while its standard deviation, theoretically expected to remain stable around 0.2, declines only slightly—likely due to the larger correlations between randomly generated universes when they are larger.
 We can see that the mean Sharpe ratio is only about 0.2 when $N=1$, the one dimensional case, which is of the same order of magnitude as the noise level.
-This highlights how diversification and data aggregation—through projection onto the ARPARP—help reduce the noise-to-signal ratio of the measurements.
+This highlights how diversification and data aggregation—through projection onto the ARPARP—help reduce the noise-to-signal ratio of the measurements when $N$ is large.
 We can also see that the exact standard deviation at $N=1$ of 0.2087 is not significantly larger than the standard deviation of 0.175, just explained by the noise of Sharpe ratio using 32.58 years of data, that we would also obtain with an infinite number of random trials instead of 20, assuming that the autocorrelation coefficients were universal across asset classes.
-This last observation justifies our choice to aggregate data across asset classes.
+This last observation, which does not conflict with the assumption of universal autocorrelation parameters across different sectors, justifies our choice to aggregate data across asset classes.
 Figure 6: Empirical Sharpe ratio based on the whole period 1990-2023 for trend following strategies using $\mathrm{ARP}$ (Eq.13 ) applied to $\mathrm{EMA}$ with $\eta=1/120$ as signal v.s.
@@
 6, our version of the renormalized $\mathrm{EMA}$, their results challenge the assumption of a purely gaussian diffusive return process from Eq.
-1, which underpins the theoretical optimality of $\mathrm{EMA}$-based strategies.
+1, which underpins the theoretical optimality of $\mathrm{EMA}$-based strategies (i.e., we would obtain $\beta$ from Eq.
+1 not as a constant, but as a path-dependent quantity that depends on $\sum_{k=1}^{t-1}(1-\lambda)^{t-1-k}\xi_{i,k}$.).
 Interestingly, when applying the Agnostic Risk Parity (ARP) strategy using $\phi-c\phi^{3}$ instead of $\phi$ as the trend signal, we observe no improvement in the Sharpe ratio.

### v7

@@
 As a consequence, using a complex basket of different complex indicators as signal, do not seem to be so rational or optimal and exposes to the risk of cherry-picking.
-November 3, 2025
+November 18, 2025
 I thank Professors Campbell Harvey, Robert Kosowsky, Valeriy Zakamulin, Denis Grebenkov and Raphael Douady for their feedbacks and valuable insights.
@@
 For example, Benveniste at al (2024) explained, for the celebration of the 50th year of the Journal of Portfolio Management, the myth surrounding the alleged inadaptation of mean–variance optimization to the real world, and how efficient mean-variance optimization works without any conflict with empirical measurements, as long as the correlation matrix is properly modeled, yet neither referenced Benichou at al (2017) nor addressed the impact of uncertainty in mean within Markowitz formula.
+The literature has long acknowledged the sensitivity of the Markowitz formula to parameter uncertainty, without recognizing that this uncertainty can actually alter the form of the formula itself [^2] A dedicated review of the literature is provided in the appendix..
 In reality, one hidden assumption of Markowitz optimization is that the expected means of returns from the signals are known and certain, but this assumption is incorrect.
@@
 The linear trading rule can be extended from the one-dimensional case to the multi-dimensional setting, allowing the incorporation of the correlation matrix to better adjust portfolio positions in order to target the maximum Sharpe ratio at the portfolio level.
-Moreover, Grebenkov and Serror (2014) is the only one to have derived an elegant333the formula proposed by Acar (2003), which depends on the correlation between the indicator and returns, leads-when expressed using the Grebenkov process-to an extremely cumbersome formula linking the parameters of the return process and the relaxation time of the EMA, which can hardly be described as objectively elegant.
+Moreover, Grebenkov and Serror (2014) is the only one to have derived an elegant444the formula proposed by Acar (2003), which depends on the correlation between the indicator and returns, leads-when expressed using the Grebenkov process-to an extremely cumbersome formula linking the parameters of the return process and the relaxation time of the EMA, which can hardly be described as objectively elegant.
 Furthermore, Zakamulin and Giner (2022) and Zakamuli and Giner (2024) are both based on the formula from Acar (2003) and rigorous formula for the theoretical Sharpe ratio, explicitly linking it to the parameters of the return process and the relaxation time of the $\mathrm{EMA}$.
 That formula is key in our main objective as we want to validate it using empirical measurements.
-Two years later, while Grebenkov and Serror (2014) remained largely unrecognized by the academic finance community 444the formula in Acar (2003); Zakamulin and Giner (2022); Zakamuli and Giner (2024) holds only with binary trading rules.
+Two years later, while Grebenkov and Serror (2014) remained largely unrecognized by the academic finance community 555the formula in Acar (2003); Zakamulin and Giner (2022); Zakamuli and Giner (2024) holds only with binary trading rules.
 Following Grebenkov and Serror (2014), Ferreira et al.
@@
 We used daily returns from 70 futures instruments in stock indices, bonds, fx and commodities futures.
-The description is in the appendix A.
-We tested the different indicators applying the $\mathrm{ARP}$ formula Eq.13. The different parameters are described in Tab 3 in the appendix B.
+The description is in the appendix B.
+We tested the different indicators applying the $\mathrm{ARP}$ formula Eq.13. The different parameters are described in Tab 3 in the appendix C.
 Fig.4 displays the sensitivities of these indicators to past daily returns. $\mathrm{MACD}$ enables us to put more weights on very older returns as we expected.
@@
 Our analysis is that market behavior should have changed in the last 100 years and we believe that our fit appears more robust than a simple variogram.
-We can see based on the Tab 4 in appendix C that empirical Sharpe is $1.24$ for $\mathrm{ARP}\left(120\right)$ with one time scale and $1.18$ for $\mathrm{MACD}\left(20,120,0.4\times 400\right)$ with 3 time scales.
+We can see based on the Tab 4 in appendix D that empirical Sharpe is $1.24$ for $\mathrm{ARP}\left(120\right)$ with one time scale and $1.18$ for $\mathrm{MACD}\left(20,120,0.4\times 400\right)$ with 3 time scales.
 $\mathrm{MACD}$ could not be justified as additional time scales do not bring significant improvement.
@@
 5.5 Correlations between indicators
-Fig.8 in Appendix C shows that the indicators with different parameters give very correlated strategies which confims that the traditional approach of using a basket of many indicators as signal is not so appealing.
+Fig.8 in Appendix D shows that the indicators with different parameters give very correlated strategies which confims that the traditional approach of using a basket of many indicators as signal is not so appealing.
 $\mathrm{ARP}\left(80\right)$ is correlated to $\mathrm{ARP}\left(150\right)$ with a coefficient of $0.96$.
@@
 No funding was received
-8 Disclosure of interest
-There are no interests to declare,
-Appendix A Data
+Appendix A Review of Literature of Solving Markowitz Optimization with Uncertainty of the Parameters
+The sensitivity of the classical Markowitz (1952) mean–variance optimization framework to estimation errors in expected returns and covariances has long been recognized.
+When means and covariances are replaced by their empirical estimates, the resulting optimal portfolios often exhibit extreme positions and poor out-of-sample performance.
+This issue has motivated a wide array of research aiming to account for parameter uncertainty in portfolio selection.
+Bayesian and Empirical–Bayes Approaches
+The Bayesian framework addresses parameter uncertainty by placing prior distributions on the mean vector and, sometimes, the covariance matrix.
+Jorion (1986) introduced a Bayes–Stein shrinkage estimator in which sample means are shrunk toward a common value, yielding more stable estimates of expected returns.
+Later, Black and Litterman (1992) proposed a practical Bayesian model combining equilibrium-implied returns with subjective investor views, producing posterior expected returns that are robust to estimation noise.
+Kan and Zhou (2007) formalized the incorporation of estimation risk directly into the optimization problem, deriving analytical adjustments to account for parameter uncertainty.
+However, these Bayesian and empirical–Bayes methods only use uncertainty information to update or regularize the input parameters of the Markowitz optimization; the structure of the optimal solution itself, $w^{\star}\propto C^{-1}\mu/\Sigma$, remains unchanged.
+Covariance Shrinkage and Regularization Approaches
+A complementary literature focuses on improving the estimation of the covariance matrix.
+Ledoit and Wolf (2003, 2004) developed optimal shrinkage estimators combining the sample covariance matrix with structured targets, leading to more stable and better-conditioned inputs for portfolio optimization.
+Jagannathan and Ma (2003) showed that imposing simple portfolio constraints (such as nonnegativity) implicitly regularizes the covariance matrix, mitigating estimation error effects.
+These approaches improve empirical robustness but, again, modify only the inputs of the optimization, not the analytical form of the optimal portfolio weights.
+Related structured-ensemble, predict-then-optimize methods steer diversification ex ante by tuning diversity in prediction and selection, using the ensemble’s combiner (equal-weight under MSE) as the target, while leaving the Markowitz closed form intact Rodriguez Dominguez at al (2025).
+Robust Optimization and Worst-Case Formulations
+Another important strand of research uses robust optimization to address parameter uncertainty.
+Goldfarb and Iyengar (2003) and El Ghaoui et al.
+(2003) recast the mean–variance problem as a convex program that optimizes for the worst-case scenario within an uncertainty set for the parameters.
+Robust optimization offers clear protection against misspecification of $\mu$ and $C$, but the resulting portfolios correspond to a modified optimization problem rather than a redefinition of the Markowitz formula itself.
+The uncertainty is modeled exogenously, not endogenously reflected in the structure of the optimal solution.
+The robust mean variance objective function is then typically
+$\min_{w}\;w^{\top}(C)w-\lambda\,\hat{\mu}^{\top}w+\lambda\rho\,\|Q^{1/2}w\|_{2}$
+It can be decomposed into three distinct components, each corresponding to a different source of risk or model uncertainty:
+  1. 1.
+    Covariance risk term: The quadratic form $w^{\top}(C)w$ represents the classical portfolio variance.
+  2. 2.
+    Expected return term: The linear component $\lambda\,\hat{\mu}^{\top}w$ corresponds to the standard mean–variance trade-off.
+The coefficient $\lambda$ controls the investor’s risk aversion and governs the balance between expected return and variance minimization.
+  3. 3.
+    Uncertainty penalization term: The additional penalty $\lambda\rho\,\|Q^{1/2}w\|_{2}$ arises directly from the ellipsoidal uncertainty set on the mean vector.
+So $\lambda\,\hat{\mu}^{\top}w-\lambda\rho\,\|Q^{1/2}w\|_{2}$ is the worst-case scenario when the $\mu$ are set in a conservative way.
+Intuitively, it reduces exposure to directions in which the mean estimates are most uncertain (as encoded by the covariance matrix $Q$ of estimation errors).
+The scalar $\rho$ acts as a robustness parameter: larger values of $\rho$ correspond to more conservative portfolios that perform well even under unfavorable perturbations of $\hat{\mu}$.
+In summary, the first term controls sensitivity to risk, the second captures expected performance under the nominal model, and the third explicitly regularizes against uncertainty in the mean estimates.
+Together, these terms yield a portfolio that trades off performance and robustness in a principled convex optimization framework.
+To our knowledge, this is the only approach in the literature that explicitly accounts for the covariance matrix $Q$ of estimation errors, although the results differ from those of Benichou at al (2017) and Valeyre (2024).
+Towards Uncertainty-Dependent Optimal Formulas
+While the above approaches (Except for the robust approach, which, however, neither relates directly to the optimal solution knowing uncertainties) treat uncertainty through priors, constraints, or uncertainty sets, they do not alter the analytical dependence of the optimal portfolio on $\mu$ and $C$.
+In contrast, Benichou at al (2017) and Valeyre (2024) propose models in which the optimal solution explicitly incorporates the uncertainty of parameter estimation into its structure.
+These works derive formulas which are different from the Markowitz one.
+In particular, under specific random matrix assumptions on the structure of the estimation-error correlation matrix, the optimal portfolio involves weighting proportional to $C^{-1/2}$ rather than $C^{-1}$, thus fundamentally revising the Markowitz prescription.
+This strand remains rare in the literature and represents a promising direction for integrating measurement uncertainty into the core of portfolio theory.
+Appendix B Data
 Commodities
@@
 Table 2: List of the 41 instruments among currencies, equity indices and Bonds.
-Appendix B Parameters
+Appendix C Parameters
 portfolio comment indicator type
@@
 Table 3: Different parameters fo the trend indicator
-Appendix C Results
+Appendix D Results
 portfolio indicator type Gross Sharpe ratio average holding period (days))

### v8

@@
 As a consequence, using a complex basket of different complex indicators as signal, do not seem to be so rational or optimal and exposes to the risk of cherry-picking.
-November 18, 2025
+December 4, 2025
 I thank Professors Campbell Harvey, Robert Kosowsky, Valeriy Zakamulin, Denis Grebenkov and Raphael Douady for their feedbacks and valuable insights.
@@
 Many financial researchers have proposed various methods to address this issue, which can be broadly classified into three categories, as noted by Koshiyama and Firoozye (2019): data snooping, overestimated performance, and cross-validation evaluation.
-Koshiyama and Firoozye (2019) introduce a technique to mitigate financial overfitting called the Covariance-Penalty Correction, which adjusts a risk metric downward based on the number of parameters and the volume of data underpinning a trading strategy.
+They introduced a technique to mitigate financial overfitting called the Covariance-Penalty Correction, which adjusts a risk metric downward based on the number of parameters and the volume of data underpinning a trading strategy.
 An alternative is to model the returns of the underlying assets using stochastic processes.
@@
 In reality, one hidden assumption of Markowitz optimization is that the expected means of returns from the signals are known and certain, but this assumption is incorrect.
-Valeyre (2024) proved that the Agnostic Risk Parity ($\mathrm{ARP}$) approach was optimal in the context of trend following, where expected means are not certain, are estimated with some errors and do not break rotational invariance as biases introduced by risk premia would, when the correlation matrix between the trends, signals or uncertainty of mean was a kind of random matrix where a dominant factor of this correlation matrix was very dominant and diffused randomly.
+Valeyre (2024) proved that the Agnostic Risk Parity ($\mathrm{ARP}$) approach was optimal in the context of trend following, where expected means are not certain, are estimated with some errors and do not break rotational invariance as biases introduced by risk premia would.
+This optimality holds when the correlation matrix between the trends, signals or uncertainty of mean is different from the correlation matrix between returns but is a kind of random matrix where a dominant factor of this correlation matrix is very dominant and diffuses randomly.
 Valeyre (2024) suggested an interacting agents model to justify such an hypothesis.
@@
 The linear trading rule can be extended from the one-dimensional case to the multi-dimensional setting, allowing the incorporation of the correlation matrix to better adjust portfolio positions in order to target the maximum Sharpe ratio at the portfolio level.
-Moreover, Grebenkov and Serror (2014) is the only one to have derived an elegant444the formula proposed by Acar (2003), which depends on the correlation between the indicator and returns, leads-when expressed using the Grebenkov process-to an extremely cumbersome formula linking the parameters of the return process and the relaxation time of the EMA, which can hardly be described as objectively elegant.
+Moreover, Grebenkov and Serror (2014) is the only one to have derived an elegant444the formula proposed by Acar (2003), which depends on the correlation between the indicator and returns, leads-when expressed using the Grebenkov process-to an extremely cumbersome formula linking the parameters of the return process and the relaxation time of the EMA, which can hardly be described as objectively elegant (description of the different formulas and the derivations when assuming the process of Grebenkov are available in the Appendix B).
 Furthermore, Zakamulin and Giner (2022) and Zakamuli and Giner (2024) are both based on the formula from Acar (2003) and rigorous formula for the theoretical Sharpe ratio, explicitly linking it to the parameters of the return process and the relaxation time of the $\mathrm{EMA}$.
 That formula is key in our main objective as we want to validate it using empirical measurements.
-Two years later, while Grebenkov and Serror (2014) remained largely unrecognized by the academic finance community 555the formula in Acar (2003); Zakamulin and Giner (2022); Zakamuli and Giner (2024) holds only with binary trading rules.
+Two years later, while Grebenkov and Serror (2014) remained largely unrecognized by the academic finance community 555the formula in Acar (2003); Zakamulin and Giner (2022); Zakamuli and Giner (2024) holds only with binary trading rules (proof in Appendix B.3, originaly derived by Firoozye and Koshiyama (2020)).
 Following Grebenkov and Serror (2014), Ferreira et al.
@@
 1 which is assumed for the underlying assets, we conducted empirical backtests of the Agnostic Risk Parity (ARP) strategy introduced by Benichou at al (2017).
-We decided to use the $ARP$ as the projection portfolio, as it is theoretically and empirically considered to be the optimal trend-following portfolio, capable of minimizing measurement noise in the Sharpe ratio which is empirically optimal (Benichou at al (2017)).
+We decided to use the $ARP$ as the projection portfolio, as it is theoretically and empirically considered to be the optimal trend-following portfolio, capable of minimizing measurement noise in the Sharpe ratio).
 However, we expect that the results should also hold even with lower Sharpe ratios and more noises for more conventional and naive ’1/N’ trend-following portfolios, such as those that do not manage correlations or assume an identity correlation matrix except when computing volatility used to target the size of the portfolio.
@@
 We used daily returns from 70 futures instruments in stock indices, bonds, fx and commodities futures.
-The description is in the appendix B.
-We tested the different indicators applying the $\mathrm{ARP}$ formula Eq.13. The different parameters are described in Tab 3 in the appendix C.
+The description is in the appendix C.
+We tested the different indicators applying the $\mathrm{ARP}$ formula Eq.13. The different parameters are described in Tab 3 in the appendix D.
 Fig.4 displays the sensitivities of these indicators to past daily returns. $\mathrm{MACD}$ enables us to put more weights on very older returns as we expected.
@@
 Our analysis is that market behavior should have changed in the last 100 years and we believe that our fit appears more robust than a simple variogram.
-We can see based on the Tab 4 in appendix D that empirical Sharpe is $1.24$ for $\mathrm{ARP}\left(120\right)$ with one time scale and $1.18$ for $\mathrm{MACD}\left(20,120,0.4\times 400\right)$ with 3 time scales.
+We can see based on the Tab 4 in appendix E that empirical Sharpe is $1.24$ for $\mathrm{ARP}\left(120\right)$ with one time scale and $1.18$ for $\mathrm{MACD}\left(20,120,0.4\times 400\right)$ with 3 time scales.
 $\mathrm{MACD}$ could not be justified as additional time scales do not bring significant improvement.
@@
 5.5 Correlations between indicators
-Fig.8 in Appendix D shows that the indicators with different parameters give very correlated strategies which confims that the traditional approach of using a basket of many indicators as signal is not so appealing.
+Fig.8 in Appendix E shows that the indicators with different parameters give very correlated strategies which confims that the traditional approach of using a basket of many indicators as signal is not so appealing.
 $\mathrm{ARP}\left(80\right)$ is correlated to $\mathrm{ARP}\left(150\right)$ with a coefficient of $0.96$.
@@
 Kan and Zhou (2007) formalized the incorporation of estimation risk directly into the optimization problem, deriving analytical adjustments to account for parameter uncertainty.
-However, these Bayesian and empirical–Bayes methods only use uncertainty information to update or regularize the input parameters of the Markowitz optimization; the structure of the optimal solution itself, $w^{\star}\propto C^{-1}\mu/\Sigma$, remains unchanged.
+However, these Bayesian and empirical–Bayes methods only use uncertainty information to update or regularize the input parameters of the Markowitz optimization; the structure of the optimal solution itself, $w^{\star}\Sigma\propto C^{-1}\mu/\Sigma$, remains unchanged.
 Covariance Shrinkage and Regularization Approaches
@@
 Together, these terms yield a portfolio that trades off performance and robustness in a principled convex optimization framework.
-To our knowledge, this is the only approach in the literature that explicitly accounts for the covariance matrix $Q$ of estimation errors, although the results differ from those of Benichou at al (2017) and Valeyre (2024).
+Segonne (2025) also uses $Q$ in a different way and introduces an anisotropy penalty into the mean-variance objective function to improve robustness.
+The term $\lambda\rho\,\|Q^{1/2}w\|_{2}$ is transformed into an anisotropy component through his Equation (48), which also helps interpret the solution of Benichou at al (2017) as a robust one.
+To our knowledge, this ”Robust Optimization and Worst-Case Formulations” is the only approach in the literature that explicitly accounts for the covariance matrix $Q$ of estimation errors, although the results differ from those of Benichou at al (2017) and Valeyre (2024).
 Towards Uncertainty-Dependent Optimal Formulas
 While the above approaches (except for the robust approach, which, however, neither relates directly to the optimal solution knowing uncertainties) treat uncertainty through priors, constraints, or uncertainty sets, they do not alter the analytical dependence of the optimal portfolio on $\mu$ and $C$.
-In contrast, Benichou at al (2017) and Valeyre (2024) propose models in which the optimal solution explicitly incorporates the uncertainty of parameter estimation into its structure.
-These works derive formulas which are different from the Markowitz one.
-In particular, under specific random matrix assumptions on the structure of the estimation-error correlation matrix, the optimal portfolio involves weighting proportional to $C^{-1/2}$ rather than $C^{-1}$, thus fundamentally revising the Markowitz prescription.
-This strand remains rare in the literature and represents a promising direction for integrating measurement uncertainty into the core of portfolio theory.
-Appendix B Data
+Brandt and Santa Clara (2006) originally introduced the idea of optimising a portfolio under the assumption that portfolio weights are proportional to time-dependent signals $x_{t}$ (not only trend-following signals).
+In their framework, this is equivalent to determining an optimal matrix $A$ such that, in their notation, $w_{t}\propto Ax_{t}$ yields the optimal ”dynamic” portfolio.
+They derived a very complex expression for $A$, although it nonetheless resembles the structure of the Markowitz solution.
+Grebenkov and Serror (2015) also optimises the portfolio by determining a matrix $A$ under the same proportionality assumption, but applies it specifically to the Grebenkov diffusive model and to exponential moving-average signals.
+Their results were derived independently from those of Brandt and Santa Clara (2006).
+Under suitable approximations, Valeyre (2024) showed that this matrix can be written as $C^{-1}C_{\xi}C^{-1}$, where $C_{\xi}$ is the covariance matrix of the estimation uncertainties $\xi$ of Eq.
+1.
+By assuming a random diffusive model for $C_{\xi}$, Valeyre (2024) obtained the ARP portfolio, for which $C^{-1}C_{\xi}C^{-1}$ reduces to $C^{-1/2}$.
+Subsequently, Firoozye et al (2023); Kelly et al (2022) proposed simplified approximations from results of Brandt and Santa Clara (2006) leading to an easy-to-use ”generalised” Markowitz formula that properly incorporates $Q$ through their Eq.
+(6), namely $C^{-1}C_{rs}Q^{-1}$ (with our notations), instead of the $C^{-1}$ appearing in the classical Markowitz formula.
+Here, $C_{rs}$ denotes the ”cross” variance matrix between returns and signals.
+The expressions $C^{-1}C_{rs}Q^{-1}$ and $C^{-1}C_{\xi}C^{-1}$ become equivalent—or at least very close—under appropriate approximations when the signals are generated both by the Grebenkov diffusive process and by exponential moving averages.
+One may therefore say that Firoozye et al (2023); Kelly et al (2022) generalised the Markowitz formula, although this contribution has remained relatively unnoticed.
+Their work also generalises Grebenkov and Serror (2015), as it is not restricted to trend-following signals nor to the Grebenkov diffusion framework.
+All these unifying relationships between the different approaches of Grebenkov and Serror (2015); Kelly et al (2022); Firoozye et al (2023); Valeyre (2024); Benichou at al (2017) are clearly explained in Segonne (2025).
+Ultimately, it appears critical to model $C_{\xi}$ (or both $C_{rs}$ and $Q$) appropriately, since these matrices are largely unknown and cannot be measured with precision without causing overfiting to the portfolio.
+Modelling them through a random diffusive process naturally leads to the solution of Benichou at al (2017), ensuring that the resulting optimal portfolio is both robust and genuinely out-of-sample optimal.
+Empirical backtests, in the context of trend-folliwng, confirm that this approach delivers strong out-of-sample performance.
+Benichou at al (2017) and Valeyre (2024) therefore propose frameworks in which the optimal portfolio explicitly incorporates estimation uncertainty into its structure.
+These works derive solutions that differ fundamentally from the classical Markowitz prescription.
+In particular, under specific random matrix assumptions on the structure of the estimation-error correlation matrix, the optimal weights become proportional to $C^{-1/2}$ rather than $C^{-1}$, thus providing a substantial conceptual revision of the Markowitz approach.
+This line of research remains rare in the literature but represents a promising direction for embedding measurement uncertainty at the core of portfolio theory.
+Appendix B Comparaison of Grebenkov’s Formula with the Litterature’s ones
+B.1 Acar (2003)
+Acar (2003) uses a binary ”Long/Short” strategy (Eq.
+17) in Section 8.5.1, ”Optimal long/short fund on the index itself XX” using his notations, where $F$ is the forecast used to predict the index $X$ and $H$ is the binary strategy.
+Acar (2003) presents his equation (Eq.
+19) in Section 8.1, based on Acar (1998).
+$H=\begin{cases}X&\text{if }F>0,\\ -X&\text{if }F<0.\end{cases}$ (17)
+$E(H_{\max})$ $=$ $\mu_{x}\left(1-2\,\phi\Big(-\frac{\mu_{x}}{\sigma_{x}\rho_{xf}}\Big)\right)+\sigma_{x}\sqrt{\frac{2}{\pi}}\rho_{xf}\exp\Big(-\frac{1}{2}\frac{\mu_{x}^{2}}{\sigma_{x}^{2}\rho_{xf}^{2}}\Big),$ (18)
+$SR(H_{\max})$ $=$ $\frac{E(H_{\max})}{\sqrt{\mu_{x}^{2}+\sigma_{x}^{2}-E(H_{\max})^{2}}}.$ (19)
+In Eq.
+19, we also need to derive $\rho_{xf}$, the correlation between the returns $X$ and the forecast $F$.
+This is not straightforward in the case of the simplest process which is used by Grebenkov and Serror (2014), to determine $\rho_{xf}$ in fonction of the paramters of the diffusive process and the parameter in the EMA used as the indicator.
+Julien Drouhet, my collegue in Machina Capital, derived the monstruous Eq.
+39 and Eq.
+40 in Sect B.2.
+We can not consider the formula of Eq.
+19 as a ’elegant’ and not obvious closed solution as the $\rho_{xf}$ need to be determined through a long derivation.
+B.2 Closed formula derived from Acar (2003) and Grebenkov and Serror (2014) process
+These computations were derived by Julien Drouhet, a collegue of mine.
+Acar’s formula
+Let $F$ be a forecast for the price of an instrument and $X$ be its future returns.
+Let $a$ and $b$ be two real numbers.
+Let’s assume that the vector $(X,F)$ is a bivariate normal distribution and let $H$ be the returns generated by the binary forecasting rule that buys $a$ units of the instrument if $F>0$ and buys $b$ units if $F<0$.
+In his article, Acar (2003) develops a closed formula that gives the expected value and variance of $H$ (hence of its Sharpe ratio) as a function of the expected value, the variance, and the correlation of $X$ and $F$ (resp.
+$\mu_{x}$, $\mu_{f}$, $\sigma_{x}^{2}$, $\sigma_{f}^{2}$ and $\rho_{xf}$):
+${\mathbf{E}}(H)$ $=\mu_{x}\left[a\Phi\!\left(\frac{\mu_{f}}{\sigma_{f}}\right)+b\Phi\!\left(\frac{-\mu_{f}}{\sigma_{f}}\right)\right]+\sigma_{x}\frac{(a-b)}{\sqrt{2\pi}}\rho_{xf}\exp\!\left(-0.5\frac{\mu_{f}^{2}}{\sigma_{f}^{2}}\right)$ (20)
+${\mathbf{E}}(H^{2})$ $=\mu_{x}^{2}\left[a^{2}\Phi\!\left(\frac{\mu_{f}}{\sigma_{f}}\right)+b^{2}\Phi\!\left(\frac{-\mu_{f}}{\sigma_{f}}\right)\right]+2\mu_{x}\sigma_{x}\frac{(a^{2}-b^{2})}{\sqrt{2\pi}}\rho_{xf}\exp\!\left(-0.5\frac{\mu_{f}^{2}}{\sigma_{f}^{2}}\right)$
+$\quad+\sigma_{x}^{2}\Bigg[a^{2}\left(\frac{\rho_{xf}^{2}}{\sqrt{2\pi}}\left(-\frac{\mu_{f}}{\sigma_{f}}\right)\exp\!\left(-0.5\frac{\mu_{f}^{2}}{\sigma_{f}^{2}}\right)+\Phi\!\left(\frac{\mu_{f}}{\sigma_{f}}\right)\right)$
+$\quad\quad+b^{2}\left(\frac{\rho_{xf}^{2}}{\sqrt{2\pi}}\left(\frac{\mu_{f}}{\sigma_{f}}\right)\exp\!\left(-0.5\frac{\mu_{f}^{2}}{\sigma_{f}^{2}}\right)+\Phi\!\left(-\frac{\mu_{f}}{\sigma_{f}}\right)\right)\Bigg]$ (21)
+In particular, when $a=1$ and b=1b=-1, the formulas read:
+${\mathbf{E}}(H)=\mu_{x}\left[1-2\Phi\left(\frac{-\mu_{x}}{\sigma_{x}\rho_{xf}}\right)\right]+\sigma_{x}\sqrt{\frac{2}{\pi}}\rho_{xf}\exp\left(-0.5\frac{\mu_{x}^{2}}{\sigma_{x}^{2}\rho_{xf}^{2}}\right)$ (22)
+$SR(H)=\frac{{\mathbf{E}}(H)}{\sqrt{\mu_{x}^{2}+\sigma_{x}^{2}-{\mathbf{E}}(H)^{2}}}$ (23)
+Grebenkov’s dynamic
+In Grebenkov and Serror (2014), the authors assume the following dynamics for the returns of the underlying price (Eq. (3), p.5 of Grebenkov and Serror (2014)):
+$r_{t}=\varepsilon_{t}+\beta\sum_{k=1}^{t-1}(1-\lambda)^{t-1-k}\xi_{t},$ (24)
+where the $\varepsilon_{k}$’s and the $\xi_{k}$’s are i.i.d following $\mathcal{N}(0,1)$.
+In matrix form, we can write the vector of the returns $\mathbf{r}$ as:
+$\mathbf{r}=\varepsilon+\beta\mathbf{E}_{1-\lambda}\xi,$ (25)
+where $\mathbf{E}_{q}$ is the $T\times T$ matrix whose coefficients $(\mathbf{E}_{q})_{j,k}$ are 0 if $j\leq k$ and $q^{j-k-1}$ otherwise.
+We directly see from the matrix expression that $\mathbf{r}$ is a centered gaussian vector, and that its covariance matrix $\mathbf{C}$ satisfies (Eq. (6), p.6 of Grebenkov and Serror (2014)):
+$\mathbf{C}=\mathbf{I}_{T}+\beta^{2}\mathbf{E}_{1-\lambda}\mathbf{E}_{1-\lambda}^{T}.$ (26)
+In particular, for two indices $j,k$, we have:
+$\mathbf{C}_{j,k}=\delta_{j,k}+\frac{\beta^{2}}{\lambda(2-\lambda)}\left[(1-\lambda)^{|j-k|}-(1-\lambda)^{j+k-2}\right].$ (27)
+Momentum signal
+Given this dynamic, Grobenkov defines a momentum signal as an EMA of the returns as follows (Eq. (13), p.8):
+$s_{t}=\gamma\sum_{k=1}^{t-1}(1-\eta)^{t-1-k}r_{k}.$ (28)
+B.2.1 From Acar’s formula to Grebenkov’s
+Covariance computation
+We have:
+$\langle s_{t},r_{t}\rangle$ $=\left\langle\gamma\sum_{k=1}^{t-1}(1-\eta)^{t-1-k}r_{k},r_{t}\right\rangle$
+$=\gamma\sum_{k=1}^{t-1}(1-\eta)^{t-1-k}\mathbf{C}_{k,t}$
+$=\gamma\sum_{k=1}^{t-1}(1-\eta)^{t-1-k}\frac{\beta^{2}}{\lambda(2-\lambda)}\left[(1-\lambda)^{|t-k|}-(1-\lambda)^{t+k-2}\right]$
+$=\gamma\frac{\beta^{2}}{\lambda(2-\lambda)}\left(\sum_{k=1}^{t-1}(1-\eta)^{t-1-k}(1-\lambda)^{t-k}-\sum_{k=1}^{t-1}(1-\eta)^{t-1-k}(1-\lambda)^{t+k-2}\right)$
+$=\gamma\frac{\beta^{2}}{\lambda(2-\lambda)}(1-\eta)^{t-1}(1-\lambda)^{t}\left(\sum_{k=1}^{t-1}(1-\eta)^{-k}(1-\lambda)^{-k}-\sum_{k=1}^{t-1}(1-\eta)^{-k}(1-\lambda)^{k-2}\right)$
+$=\gamma\frac{\beta^{2}}{\lambda(2-\lambda)}(1-\eta)^{t-1}(1-\lambda)^{t}\left(\sum_{k=1}^{t-1}\left(\frac{1}{(1-\eta)(1-\lambda)}\right)^{k}-\frac{1}{(1-\lambda)^{2}}\sum_{k=1}^{t-1}\left(\frac{1-\lambda}{1-\eta}\right)^{k}\right)$
+$=\gamma\frac{\beta^{2}}{\lambda(2-\lambda)}(1-\eta)^{t-1}(1-\lambda)^{t}\left(\frac{\frac{1}{(1-\eta)(1-\lambda)}-\left(\frac{1}{(1-\eta)(1-\lambda)}\right)^{t}}{1-\frac{1}{(1-\eta)(1-\lambda)}}-\frac{1}{(1-\lambda)^{2}}\frac{\frac{1-\lambda}{1-\eta}-\left(\frac{1-\lambda}{1-\eta}\right)^{t}}{1-\frac{1-\lambda}{1-\eta}}\right)$
+We could simplify further, but the final expression isn’t much more tractable.
+If we let $t\rightarrow\infty$, we obtain:
+$\langle s_{t},r_{t}\rangle=\gamma\frac{\beta^{2}}{\lambda(2-\lambda)}\frac{\frac{1}{1-\eta}}{\frac{1}{(1-\eta)(1-\lambda)}-1},$ (29)
+that is:
+$\langle s_{t},r_{t}\rangle=\gamma\frac{\beta^{2}}{\lambda(2-\lambda)}\frac{1-\lambda}{1-(1-\eta)(1-\lambda)}.$ (30)
+As in Grebenkov and Serror (2014), we rescale the parameter $\beta$ in order to make the asymptotic variance independent of the timescale $\lambda$ by defining:
+$\beta_{0}=\frac{\beta}{\sqrt{\lambda(2-\lambda)}},$ (31)
+so that the asymptotic variance of the returns in (26) becomes:
+$\sigma_{x,\infty}^{2}=1+\beta_{0}^{2}.$ (32)
+The covariance becomes:
+$\langle s_{\infty},r_{\infty}\rangle=\gamma\beta_{0}^{2}\frac{1-\lambda}{1-(1-\eta)(1-\lambda)}.$ (33)
+Momentum’s variance computation
+Before computing the variance of $s_{t}$, we state this lemma, that is a simple computation:
+Lemma B.1.
+Let $p$ and $q$ be two real numbers, and $N$ a positive integer. One has:
+$\sum_{1\leq l<k\leq N-1}p^{k+l}q^{k-l}=\frac{1}{1-p/q}\left(\frac{p}{q}\frac{\left(p/q\right)-\left(p/q\right)^{N}}{1-p/q}-\frac{p^{2}-p^{2N}}{1-p^{2}}\right),$ (34)
+$\sum_{1\leq l<k\leq N-1}p^{k+l}q^{k+l}=\frac{1}{1-pq}\left(pq\frac{pq-(pq)^{N}}{1-pq}-\frac{(pq)^{2}-(pq)^{2N}}{1-pq}\right).$ (35)
+We can then compute the variance $\langle s_{t},s_{t}\rangle$:
+Proposition B.1.
+Under the previous model, we have:
+$\langle s_{t},s_{t}\rangle=\gamma^{2}p^{2t-2}\left(A_{t}+2\beta_{0}^{2}\left(B_{t}+(1-\lambda)^{-2}C_{t}\right)\right),$ (36)
+where:
+$A_{t}$ $=\left(1+\beta_{0}^{2}\right)\frac{(1-\eta)^{-2}-(1-\eta)^{-2t}}{1-(1-\eta)^{-2}}-\beta_{0}^{2}(1-\lambda)^{-2}\frac{\left(\frac{1-\lambda}{1-\eta}\right)^{2}-\left(\frac{1-\lambda}{1-\eta}\right)^{2t}}{1-\left(\frac{1-\lambda}{1-\eta}\right)^{2}}$
+$B_{t}$ $=\frac{1}{1-((1-\eta)(1-\lambda))^{-1}}\left(\frac{((1-\eta)(1-\lambda))^{-2}-((1-\eta)(1-\lambda))^{t+1}}{1-((1-\eta)(1-\lambda))^{-1}}-\frac{(1-\eta)^{-2}-(1-\eta)^{-2t}}{1-(1-\eta)^{-2}}\right)$
+$C_{t}$ $=\frac{1}{1-\frac{1-\lambda}{1-\eta}}\left(\frac{1-\lambda}{1-\eta}\frac{\frac{1-\lambda}{1-\eta}-\left(\frac{1-\lambda}{1-\eta}\right)^{t}}{1-\frac{1-\lambda}{1-\eta}}-\frac{\left(\frac{1-\lambda}{1-\eta}\right)^{2}-\left(\frac{1-\lambda}{1-\eta}\right)^{2t}}{1-\left(\frac{1-\lambda}{1-\eta}\right)^{2}}\right).$
+In particular, in asymptotic regime:
+$\sigma_{f,\infty}^{2}:=\langle s_{\infty},s_{\infty}\rangle=\gamma^{2}\frac{1+\frac{\beta_{0}^{2}}{1-(1-\eta)(1-\lambda)}}{1-(1-\eta)^{2}}.$ (37)
+Proof.
+Let $p=1-\eta$ and $q=1-\lambda$. We have using 26:
+$\langle s_{t},s_{t}\rangle$ $=\gamma^{2}\sum_{1\leq k,l\leq t-1}p^{2t-2-k-l}\langle r_{t},r_{t}\rangle$
+$=\gamma^{2}\left(\sum_{k=1}^{t-1}p^{2t-2-2k}\left(1+\beta_{0}^{2}\left(1-q^{2k-2}\right)\right)+2\sum_{1\leq l<k\leq t-1}p^{2t-2-k-l}\beta_{0}^{2}\left(q^{k-l}-q^{k+l-2}\right)\right)$
+$=\gamma^{2}p^{2t-2}\left(A_{t}+2\beta_{0}^{2}B_{t}+2\beta_{0}^{2}q^{-2}C_{t}\right),$
+where:
+$A_{t}$ $=\sum_{k=1}^{t-1}p^{-2k}\left(1+\beta_{0}^{2}\left(1-q^{2k-2}\right)\right),$
+$B_{t}$ $=\sum_{1\leq l<k\leq t-1}p^{-k-l}q^{k-l},$
+$C_{t}$ $=\sum_{1\leq l<k\leq t-1}p^{-k-l}q^{k-l}.$
+Applying the lemma on $\tilde{p}=1/p$ and $\tilde{q}=q$, we can compute $A_{t}$, $B_{t}$ and $C_{t}$ and we obtain the wanted formula.
+To get the asymptotic formula, we just have to notice that:
+$p^{2t-2}A_{t}$ $\underset{t\rightarrow\infty}{\longrightarrow}\frac{1+\beta_{0}^{2}}{1-p^{2}},$
+$p^{2t-2}B_{t}$ $\underset{t\rightarrow\infty}{\longrightarrow}\frac{p^{-2}}{1-(pq)^{-2}}\frac{1}{1-p^{-2}},$
+$p^{2t-2}C_{t}$ $\underset{t\rightarrow\infty}{\longrightarrow}0,$
+which concludes after simplifying the expression. ∎
+Acar’s formula
+In this section, we assume for simplicity that we are in the stationary regime ($t\rightarrow\infty$) in order to remove border effects.
+Recall the formula for the covariance between $s_{t}$ and $r_{t}$ in this regime:
+$\langle s_{\infty},r_{\infty}\rangle=\gamma\beta_{0}^{2}\frac{1-\lambda}{1-(1-\eta)(1-\lambda)}.$ (38)
+Since the returns are centered in this model, Acar’s formula in the case of a binary signal becomes (from Eq. (22)):
+$E(H)$ $=\sqrt{\frac{2}{\pi}}\sigma_{x}\rho_{fx}$
+$=\sqrt{\frac{2}{\pi}}\frac{\langle s_{\infty},r_{\infty}\rangle}{\sqrt{\langle s_{\infty},s_{\infty}\rangle}}.$
+Using the formulas for $\langle s_{\infty},r_{\infty}\rangle$ and $\langle s_{\infty},s_{\infty}\rangle$, we obtain:
+${\mathbf{E}}(H)=\sqrt{\frac{2}{\pi}}\beta_{0}^{2}\frac{(1-\lambda)\sqrt{1-(1-\eta)^{2}}}{\sqrt{\left(1-(1-\eta)(1-\lambda)\right)\left(1+\beta_{0}^{2}-(1-\eta)(1-\lambda)\right)}}.$ (39)
+We then get:
+$SR(H)=\frac{{\mathbf{E}}(H)}{\sqrt{1+\beta_{0}^{2}-{\mathbf{E}}(H)^{2}}}.$ (40)
+B.3 Acar’s formula for linear signals
+Let’s mention a useful lemma for what follows:
+Lemma B.2.
+Let $Z\sim\mathcal{N}(0_{N},I_{N})$ be a gaussian vector. Let $A\in\mathcal{S}_{N}({\mathbb{R}})$ be an $N\times N$ positive-semidefinite matrix. Then one has:
+${\mathbf{E}}\left[Z^{T}AZ\right]=\mathrm{tr}(A),$ (41)
+${\mathbf{E}}\left[\left(Z^{T}AZ\right)^{2}\right]=2\mathrm{tr}(A^{2}).$ (42)
+Proof.
+If we diagonalize $A$, we can see that:
+$Z^{T}AZ=\sum_{i=1}^{N}\lambda_{i}P_{i}^{2},$ (43)
+where $P_{1},\cdots,P_{N}$ are i.i.d. $\mathcal{N}(0,1)$ variables, and $\lambda_{1},\cdots,\lambda_{N}$ are $A$’s (non-negative) eigenvalues.
+A direct computation gives the result. ∎
+We recall that $F$ is the forecast for the price of an instrument and $X$ is its future returns [^13] We keep the notations $\mu_{x},\mu_{f},\sigma_{x},\sigma_{f}$ and $\rho_{xf}$.
+The following result was proven by Firoozye and Koshiyama (2020).
+Proposition B.2.
+Let us assume that $(X,F)$ is a centered bivariate normal distribution and let $H$ be the returns of the (linear) strategy that trades $F$, that is:
+$H=FX.$ (44)
+Then the expected value and the Sharpe ratio of $H$ are computable:
+${\mathbf{E}}(H)=\rho_{xf}\sigma_{x}\sigma_{f}$ (45)
+$SR(H)=\frac{\rho_{xf}}{\sqrt{1+\rho_{xf}^{2}}}.$ (46)
+Proof.
+For the computation of ${\mathbf{E}}H$, we don’t assume that $(X,F)$ is centered.
+Let’s define, with obvious notations, $Z:=\left(X,F\right)^{T}\sim\mathcal{N}(\mu,\Sigma)$. Then the random vector:
+$\tilde{Z}:=\Sigma^{-1/2}(Z-\mu),$ (47)
+follows $\mathcal{N}(0_{N},I_{N})$.
+Then, if we define the matrix:
+$A=\begin{pmatrix}0&\tfrac{1}{2}\\[6.0pt] \tfrac{1}{2}&0\end{pmatrix}$, (48)
+we can see that:
+$H=Z^{T}AZ.$ (49)
+Since $Z=\mu+\Sigma^{1/2}\tilde{Z}$, we can see that:
+$H=\mu^{T}A\mu+2\mu^{T}A\Sigma^{1/2}\tilde{Z}+\tilde{Z}^{T}\Sigma^{1/2}A\Sigma^{1/2}\tilde{Z}$. (50)
+Taking the expected value of this quantity, we get:
+${\mathbf{E}}H$ $={\mathbf{E}}\left[\mu^{T}A\mu+2\mu^{T}A\Sigma^{1/2}\tilde{Z}+\tilde{Z}^{T}\Sigma^{1/2}A\Sigma^{1/2}\tilde{Z}\right$]
+$=\mu^{T}A\mu+0+{\mathbf{E}}\left[\tilde{Z}^{T}\Sigma^{1/2}A\Sigma^{1/2}\tilde{Z}\right$]
+$=\mu_{x}\mu_{f}+\mathrm{tr}\left(\Sigma^{1/2}A\Sigma^{1/2}\right)$
+$=\mu_{x}\mu_{f}+\mathrm{tr}\left(A\Sigma\right)$
+$=\mu_{x}\mu_{f}+\rho\sigma_{x}\sigma_{f}$
+For the second formula, let’s assume the $\mu=0_{2}$ and compute ${\mathbf{E}}[H^{2}]$. In this case, using the lemma:
+${\mathbf{E}}[H^{2}]$ $={\mathbf{E}}\left[(\tilde{Z}^{T}\Sigma^{1/2}A\Sigma^{1/2}\tilde{Z})^{2}\right]$
+$=2\mathrm{tr}\left(\left(\Sigma^{1/2}A\Sigma^{1/2}\right)^{2}\right)$
+$=2\mathrm{tr}\left(\Sigma^{1/2}A\Sigma^{1/2}\Sigma^{1/2}A\Sigma^{1/2}\right)$
+$=2\mathrm{tr}\left(A^{2}\Sigma^{2}\right)$
+$=\frac{1}{2}\left(\sigma_{x}^{4}+2\rho_{xf}^{2}\sigma_{x}^{2}\sigma_{f}^{2}+\sigma_{f}^{4}\right),$
+which gives the result.
+Corollary B.1.
+In Grebenkov and Serror (2014) stationary model, we have the following formulas:
+$E(H)=\langle s_{\infty},r_{\infty}\rangle,$ (51)
+$SR(H)=\frac{1}{\sqrt{1+\frac{\langle s_{\infty},s_{\infty}\rangle\langle r_{\infty},r_{\infty}\rangle}{\langle s_{\infty},r_{\infty}\rangle^{2}}}},$ (52)
+where $\langle s_{\infty},r_{\infty}\rangle$, $\langle r_{\infty},r_{\infty}\rangle$ and $\langle s_{\infty},s_{\infty}\rangle$ have closed-form formulas only depending on the model’s parameters.
+B.4 Ferreira et al. (2018)
+Ferreira et al. (2018), which cited Grebenkov and Serror (2014), used also a linear trading rule (Eq. 53) in their formula (3) from Section 2.2.2, ”Risk and Returns”.
+$\langle R\rangle=\frac{1}{T-N+1}\sum_{t=N}^{T}m_{t-1}(N)\,X_{t}$ (53)
+In Eq.
+53, $m$ is a simple moving average (SMA) of returns rather than an exponential moving average (EMA).
+Ferreira et al.
+(2018) derived the Sharpe ratio over two pages without using the results of Acar (2003), obtaining an explicit formula in contrast to Acar (2003).
+In their formula (Eq.
+54), which appears in Section 2.3, ”Limits and Interpretations”, the Sharpe ratio depends only on $\rho(t,t-i)$, the correlation coefficient of returns at times $t$ and tit-i.
+Ferreira et al.
+(2018) provided the formula closest to Grebenkov and Serror (2014).
+The formula is applicable to a more generalized diffusive process as in Grebenkov and Serror (2014), but it is less straightforward and not entirely consistent, since SMA differs from EMA.
+Ferreira et al.
+(2018) was published later and includes more parameters in the formula.
+$SR=\frac{\sum_{i=1}^{N}\rho(t,t-1)}{\sqrt{N+\Big(\sum_{i=1}^{N}\rho(t,t-i)\Big)^{2}+\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{N}\rho(t-j,t-i)}}$ (54)
+B.5 Zakamulin and Giner (2022)
+Zakamulin and Giner (2022) is also based on a simple moving average (SMA) of returns, Eq 55 (”2.
+Time Series Momentum Strategies”) but employs a binary strategy in eq 56.
+They derive the Sharpe ratio in Eq 59 using the same approach as Acar (2003), which requires the correlation between returns and the forecast ($\rho_{n}$ is ”the correlation coefficient between $r_{t}$ and the $MOM_{t-1}(n)$ ).
+This correlation is not straightforward to compute to claim explicit formula.
+$\text{MOM}_{t-1}(n)=\sum_{i=1}^{n}X_{t-i}$ (55)
+$R_{t}^{LS}=\begin{cases}r_{t},&\text{if }\text{MOM}_{t-1}(n)>0,\\[5.69054pt] 2r_{f,t}-r_{t},&\text{otherwise.}\end{cases}$ (56)
+$E[R_{t}^{LS}]$ $=$ $(2\,\phi(-d)-1)\,\mu+2\,\big(g+\phi(d)\,r_{f}\big),$ (57)
+$\text{Var}[R_{t}^{LS}]$ $=$ $(\mu^{2}+\sigma^{2})+4r_{f}\big(g-(\mu-r_{f})\phi(d)\big)-\big(E[R_{t}^{LS}]\big)^{2},$ (58)
+$g$ $=$ $\sigma\,\rho_{n}\,\phi(d)$ (59)
+Appendix C Data
 Commodities
@@
 Table 2: List of the 41 instruments among currencies, equity indices and Bonds.
-Appendix C Parameters
+Appendix D Parameters
 portfolio comment indicator type
@@
 Table 3: Different parameters fo the trend indicator
-Appendix D Results
+Appendix E Results
 portfolio indicator type Gross Sharpe ratio average holding period (days))